A Complete Bibliography of Publications in Reviews of Modern Physics (1940–1949)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

02 November 2023 Version 1.07

Title word cross-reference

2 [AB44, Ger49, Her42a]. _{2}^{+} [Mro42]. F [DC48, DC49]. γ [Lat47]. h [DC48, DC49]. k ln n! [Ste49b]. l [Her42b]. m [DC48, DC49]. μ [TW49a]. μ^{−} [Whe49]. N [DC48, DC49].

-Meson [TW49a]. -Mesons [Whe49]. -Radiation [Lat47]. -Rays [Lat47]. -Type [Her42b].

1941 [Bir41].

60th [Pau45].

70-tem [vL49].

above [Rey49]. Absolute [Dus45]. Absorber [WF45]. Absorption [Bro42, CPS42, Cha45, Cha49b, CD48, EE48, KM42, MM42, NTD44, NN44, RAO^{+}49, She42a, She42b, Skl42, Spo42, Spo44]. Abundance [Bow48]. Abundances [Bro49b]. Accelerators [Sla48]. Acids [PW48]. Actinide [Ham48]. Action [WF49]. Activated [Pri42]. Activity [Fra45]. Addendum [Bri46b]. Additional [ES46]. Ag [GI48]. Age [Ano49a]. Aggregation [She42a, She42b]. Air [Coc49]. Al [Cha49b]. Albert [vL49, Mil49a]. Alkali [Sei46]. Alkyl [Mul42c]. Altitudes [AAC49, BMN^{+}48a, BMN^{+}48b]. Analogy [Dir45]. Analysis [Hal47, MR42]. Andrews [Eps48]. Anelastic [Eck48b]. Angular [GH45]. Anisotropic [Hea46]. Annihilation [Mil49b]. Apparatus [MB42, Rey49]. Application [Dus45, Lem49, Mei42, Mor46, PW48, Wil45]. Applications [GG46]. Applied [Fis42]. Approach [BD48, Fey48]. Approximate [Eck48a]. Approximation [Wil45]. Associated [FGV49]. Astronomy [Cha43]. Atmosphere [Cha45, Whi42, Whi43, Wil42a]. Atmospheres [Cha44]. Atmospheric [MFDA49]. Atom [Mil49b]. Atom-Annihilation [Mil49b]. Atomic [Bri45b, DC48, DC49, Fis42, MB42, Mar47, Meg42a, Mei42, Smy45, Wil42b]. Atoms [BR45, KM46, PE48]. Attempt [Mei45]. Attempts [Wil42b]. Auditory [Fle40]. August [Bir41]. Average [MBH47].

Bands [McK42, Swi42, TF42]. Basis [Bha49]. Beam [Est46, Fis42, KM46, MB42, Ske46, Wil42b]. Beams [Mei42]. Benzene [Skl42, TF42]. Benzenes [Spo42]. Beta [Cra48, Kon43, MBH47]. Beta-Decay [Cra48, Kon43]. Beta-Rays [MBH47]. Between [Dir45, Bar49, Bar48, Whe49]. Biology [Loo40]. Birthday [Mil49a, Pau45]. Bodies [KSW49]. Body [Fok49]. Bohr [Pau45]. Bomb [Bri45b]. Bond [MR42]. Borderland [Loo40]. Born [Wil45]. Boundary [LB49]. Broadened [VW45]. Brownian [Cha49a, WU45]. Bursts [Cla49b]. Butadiene [Mul42c].

Caesium [KM42]. Calculation [MM42]. California [DuB48]. Canonical [BB49]. Carbon [MR42]. Case [PH45]. Catalysis [Bee45, Bee48]. Catalysts [Bee45]. Cavitating [PS48a]. Cellular [Hev45]. Centered [Lei48]. Centers [Sei46]. CH [Her42a]. Chain [Ano47, The46]. Challenge [Bee45]. Chamber [Cha49b, GG46]. Characteristics [Ano47, The46]. Charge [TW49a]. Charge-Exchange [TW49a]. Charged [HTW49]. Chemical [Kau42]. Chloride [NN44]. Chlorophyll [Rab44]. Chloroplasts [Fra45]. Clarification [Bar48]. Classical [Dir45, WF49, Wil45]. Cloud [Cha49b, GG46]. Cloud-Chamber [Cha49b]. CO [Mro42]. Coincidence [Mit48]. Collision [Mar47, VW45, Wil45]. Collision-Broadened [VW45]. Collodial [GI48]. Color [HS42, Sei46]. Cometary [McK42, Swi42]. Comets [Bob42, Her42a, Van42]. Comments [Hul45]. Compilation [GIF47]. Complications [Hel42]. Component [BFL^{+}49]. Compounds [PE48]. Concept [Bro45, vKL49]. Concepts [Bec45, Wil45]. Concerning [Coc49].

Concluding [Opp49]. Conduction [MB40]. Conference [Mul42b]. Conformal [Sch49]. Connection [Bar49]. Consequences [Whe49]. Constant [Lon45]. Constants [Bir41, DC48, DC49, Hea46]. Constitution [HS42, Wil42a]. Construction [Wil42c]. Continuous [Deu48]. Contraction [Rac49, Swa41b]. Cooperative [Wan45]. Corpuscules [dB49]. Corrections [ES46]. Cosmic [AAL^{+}48, AAC49, BMN^{+}48a, BMN^{+}48b, BFL^{+}49, Cla49a, FGV49, HW45, KSW49, Kor48, Kor49, NR48, Opp49, Roc49, RG41, Ros48, Ros49b, Val49]. Cosmic-Radiation [Cla49a]. Cosmic-Ray [BMN^{+}48a, BMN^{+}48b, HW45, KSW49, NR48, Opp49, Roc49, RG41, Ros48]. Cosmogony [Gam49]. Cosmological [Göd49, Lem49]. Council [Jew48]. Counters [Kor42]. Counting [Mit48]. Coupling [PH45]. Covalent [Sam46]. Creep [Dus45]. Critique [Kor49]. Cross [GIF47]. Crystal [Hou48, Lei48]. Crystals [Fre42, Sei46]. Cubic [Lei48]. Current [Fie46]. Cyclotrons [Liv46].

Data [BW47, Bob42]. Decay [Ano47, Cra48, Kon43, The46, TW49b]. Decomposition [Rac49]. Degenerate [Sha44]. d’Einstein [dB49]. Dependent [PH45]. Depth [Hen49]. Derivable [DC48, DC49]. Derivatives [Mul42c, Skl42]. Design [Sla48]. Detection [Rey49]. Determining [Mit48, PW48]. Deuterium [Hal47]. Deutero [NN44]. Deutero-Methyl [NN44]. Deuteron [GH45]. Developments [Wei49]. Devices [Ske46]. Diamagnetism [Bar48]. Dichroism [Hel42]. Dielectric [Kau42]. Dielectrics [MB40]. Dienes [CPS42]. Difference [Ste47]. Differences [Mei45]. Difficulties [Hul45]. Diffracting [PB42]. Diffraction [Gin43]. Diffusion [JF46]. Dilute [Hel42]. Dimensional [Eck48a]. Dipole [PW48]. Dirac [Pau43, Tau49]. Direct [WF49]. Disagree [Fra41]. Discharge [Loe48]. Discharges [DP40, DP41]. Discussion [Ano42a, Ano42e, Ano42f, RAO^{+}49, Ano42b, Ano42c, Ano42d, Ano49b]. Disintegration [GH45, Mit48, RAO^{+}49]. Dispersed [Hel42]. Disruptions [Kor48, Kor49]. Dissociation [Sam46]. Distribution [Gou45, GH45]. Division [Hev45]. Do [Fra41]. Domain [Bro45]. Domains [Kit49]. Double [Hel42]. Doubling [Her42b]. Down [Mar47]. Drag [PS48a]. Dualité [dB49]. Dyes [Bro42, HS42, She42a, She42b]. Dynamical [Cha49a]. Dynamics [Cha49a, Dir49].

Earth [Meg42a, Whi42, Whi43]. Echelons [Wil42c]. Effect [BMN^{+}48b, BW45, Cle47, Glo43, Hev45, Rey49, Val49]. Effects [Bri45a, NR48, Rab42, She42a, She42b]. Einstein [Fra49, Göd49, Mil49a, Str49]. Einsteins [vL49]. Elastic [Hea46, Mar47]. Electret [Gut48]. Electric [Hel42]. Electrical [DP40, DP41, MB40]. Electrodynamics [Hof49, PS48b, WF49]. Electromagnetic [Eli47, Whe49]. Electron [Fie46, Mor46, Rab42, SY49, Wei49]. Electronic [Mul42a, Ske46, Skl42, ST41]. Electronics [Sla46]. Electrons [Eli47, PE48, Ros49b, TW49b]. Elementary [Bha45, Bha49, Bor49, NW49, Pau41]. Elements [Gin43, GIF47, Ham48, Mar41, Meg42a]. Emission [HN49, McK42, Mro42]. Emitted [MBH47]. Emitting [SY49]. Empirical [Fee47]. Emulsions [Sha41]. Energies [PE49]. Energy [AY48, Cla49a, Cla49b, Cra48, Fee47, FL49, HL48, MBH47, Mei45, Rey49, Smy45, TW49b]. Entropy [Ste49b, TF48]. Environment [She42a, She42b]. Equations [BB49, Bha45, Eck48a, Göd49, Tau49]. Equilibrium [BD48, Kle49]. Errata [DP41, She42b]. Erratum [Ano47, SH50, Smy48a]. Especially [MB42]. Ethylene [Bee45, Mul42c]. Evidence [BFL^{+}49, Her42a, LR49, Mil49b]. Example [Göd49]. Exchange [TW49a]. Excitation [FL49, Wil42b]. Excited [FLL48]. Exemplified [Bee45]. Existence [LR49]. Expansion [ES45, ES46]. Experiment [Swa41a]. Explosion [AY48]. Extensive [Coc49, MFDA49].

F [Ger49]. Face [Lei48]. Face-Centered [Lei48]. Factors [Loe48]. Fatty [PW48]. Fe [Cha49b]. Feet [AAL^{+}48]. Ferromagnetic [Con45, Kit49]. Ferromagnetism [Van45]. Field [Bec45, BB49, Bri46a, Bri46b, Eli47, Göd49, Pau41, Pau43, Ros49a, Str49]. Fields [BR45, ES45, ES46, Fre42, Kle49, Sch49]. Film [dHvdB49]. Fission [Ano47, Ham48, Mei45, The46, Tur40]. Fitting [BW47]. Fitzgerald [Swa41b]. Five [KSW49]. Flares [FGV49, NR48]. Flow [LB49, PS48a]. Fluid [Eck48b]. Fluids [Kle49]. Fluorescence [Pri42, Win44]. Foils [Cha49b]. Forbidden [Mro44]. Forces [EES49]. Foreword [Dar44, Dar45]. Formation [Cha45]. Formed [Ano47, The46]. Forms [Dir49]. Fractionation [BD48]. Fracture [Zen45]. Freedom [Wea48]. Friction [Bik44a, Bik44b, Cha49a]. Functions [Her49, Kar44]. Fundamental [Bar48].

Galactic [Val49]. Galvanometer [Str48]. Gamma [EE48, FLL48]. Gamma-Radiation [FLL48]. Gamma-Ray [EE48]. Gaps [Smy48a, Smy48b]. Gases [DP40, DP41, MB42, NTD44]. Geburtstag [vL49]. General [Bir41, IS49]. Generalized [Ein48, PS48b]. Genesis [Jew48]. Geochemistry [Wil42a]. Geometry [Fok49]. Ghosts [PB42]. Giant [GK45]. Gold [KPC45]. Gold-Iron [KPC45]. Gold-Nickel [KPC45]. Grating [AB44, PB42]. Gravitation [ES45, ES46, Ein48, Göd49]. Gravitational [Kle49, Ros49a]. Guns [Fie46, Mor46].

Halide [Pri42, Sei46]. Heat [Lei48, MR42]. Heavenly [KSW49]. Heavy [BFL^{+}49, LR49, Ste49a, Tur45]. Helium [Dar40]. High [AAC49, BMN^{+}48a, BMN^{+}48b, Bri45a, Bri46a, Bri46b, Cla49b, Fie46]. Hydrodynamic [EES49]. Hydrogen [Cha44, MB42, Ste47, Wil42b]. Hydrogenation [Bee45]. Hydrostatic [Bri45a]. Hyperbolic [Kar44]. Hypothesis [Mil49b].

Identification [PB42]. Identifications [Swi42]. II [BB49, Den40, WU45]. Impurities [Con45]. Incidence [AB44]. Increases [FGV49]. Individual [ES45, ES46]. Induced [LS40]. Influence [ES45, ES46, KSW49, Rol48]. Infra [AB44, Bar42, Den40, NTD44, NN44, ORA44, OR44]. Infra-Red [AB44, Bar42, Den40, NTD44, NN44, ORA44, OR44]. Institute [DuB48]. Integrals [MR42]. Intensity [McK42]. Inter [FL49]. Inter-Molecular [FL49]. Interaction [BT49, Eli47, Lat47, WF45, Whe49]. Interactions [PH45]. Intermetallic [PE48]. Interparticle [WF49]. Interpretation [Ros48, Str44]. Intra [FL49]. Intra- [FL49]. Introduction [Mul42b]. Invariance [Tho45]. Invariant [PV49]. Ion [Liv46, Mro42]. Ionization [Ger49, Kor49]. Ionospheric [Ger49]. Ions [Cha44, Fre42]. Iron [KPC45]. Irreversible [TF48]. Isolated [Fra45]. Isomeric [ORA44, OR44]. Isomerism [SH49, SH50]. Isotopes [JF46, MBH47, Sea44, SP48, Ste49a]. Isotropic [vKL49].

Knowledge [DC48, DC49]. Kron [Hof49]. Kynch [MR42].

Laboratory [Mro44]. Lagrangian [Lan49]. Laminar [LB49]. Latitude [BMN^{+}48b]. Lattice [Hou48]. Laws [NTD44]. Layer [LB49]. Lead [Rey49]. Least [BW47]. Least-Square [BW47]. Length [Meg42b]. Levels [HL48]. Light [Bob42, CL48, Elv42, Fis42, FLL48, HL48]. Line [Fis42, AB44]. Linear [BB49, Her42b, Sla48]. Lines [Cha45, CD48, Gou45, Mro44, VW45]. Liquid [Gin43, Hen49, ORA44]. Localized [NW49]. L’Oeuvre [dB49]. Lorentz [Swa41b]. Low [DP40, DP41, Lon45, Rey49, Ske46].

Magnesium [Fis42]. Magnetic [BR45, BW45, KM46, NR48]. Magnetism [Boz47]. Magnetization [KPC45]. Mass [Bro49a, Ste47]. Masses [Ste49a]. Materials [Hea46]. Matter [Lat47]. Means [BW47]. Measurements [AB44, Kor49, McK42]. Mechanics [Dir45, Fey48]. Mechanism [DP40, FGV49, WF45, DP41]. Mechanisms [Loe48]. Merits [Fra41]. Meson [Cha49b, Hei49, Hul45, LR49, Sch49, TW49a, TW49b, Wen47, Yuk49]. Mesons [HW45, Rey49, RAO^{+}49, SY49, Whe49]. Mesotron [Bro49a]. Metabolic [Ham48]. Metal [Bee45, GI48]. Metals [Bri45a, Con45, Dus45, PE48]. Metaphysical [Mar41]. Meteors [Whi42, Whi43]. Method [KM46, Pau43, PW48, Tau49]. Methods [Mit48, Yuk49]. Methyl [NN44]. Micro [Con42]. Micro-Wave [Con42]. Microscopic [Cas45]. Microwave [Gor48, Sla46, Smy48a, Smy48b]. Migration [FL49]. Military [Smy45]. Millikan [DuB48, Eps48, Jew48, Rol48]. Missing [Tur45]. Model [GK45]. Models [Yuk49]. Modes [Sha44]. Molecular [Bar48, Est46, FL49, Her49, KM46, Swi42]. Molecules [Bar42, Den40, Her42a, Her42b, KM46, Mul42a, PK44, Sam46, Sha44, ST41]. Moment [PW48]. Momentum [Cra48]. Monatomic [DG41, Wil42b].

Monoergic [HTW49]. Motion [Cha49a, IS49, WU45]. Motions [Cle47]. Moving [Cha45, Fok49, Str48]. Multiplier [Lan49].

N [AB44]. Narrow [Smy48b, Smy48a]. National [Jew48]. Near [Spo44, TF42]. Needle [Str48]. Negative [Cha44]. Neutrino [Cra48]. Neutron [GIF47, Ste47]. Neutron-Hydrogen [Ste47]. Neutrons [GH45, HTW49, Mar47]. Nickel [KPC45]. Niels [Pau45]. Night [Elv42]. Nocturnal [Ger49]. Nodal [Her49]. Non [BB49, Fey48, Hof49]. Non-Linear [BB49]. Non-Relativistic [Fey48]. Non-Riemannian [Hof49]. Normal [Hou48]. Note [Bar48, PE49]. Novae [Zwi40]. Nuclear [Bro49b, Fee47, Kor48, Kor49, Mit48, RAO^{+}49, SH49, Sha41, Tur40, SH50]. Nuclei [BFL^{+}49, CL48, FLL48, HL48, Mar47, The46, Tur45, Whe49, Ano47]. Nucleus [TW49a].

O [AB44]. Oblique [AB44]. Observation [Rob49]. Observations [AAC49]. Obtaining [Wil42b]. Octanes [ORA44, OR44]. Often [Fra41]. Ondes [dB49]. One [Eck48a]. One-Dimensional [Eck48a]. Onsager [Cas45]. Operation [Kor42]. Orders [PB42]. Ordinary [Wil45]. Organic [PK44]. Origin [Cha46, Hel42, Val49]. Other [DC48, DC49, Ske46]. Our [DC48, DC49, ES46]. Oxide [Mul42a]. Oxides [GI48]. Oxygen [Bow48, Cha44].

Pair [PH45]. Paper [ES46, MR42]. Parastatic [Str48]. Part [Den40]. Partial [Win44]. Particle [HTW49, Ros49a]. Particles [Bha45, Bha49, Bor49, Cla49a, Cla49b, IS49, Pau41, Roc49, Sha41]. Partition [AY48]. Patterns [Fle40]. Pb [Cha49b]. Penetrating [MFDA49, Roc49]. Penney [MR42]. Perturbations [Sha44]. Phase [Bar49, ORA44, OR44]. Phenomena [AY48, MB40, Ros48, Wan45]. Philosophers [Fra41]. Philosophy [Fra49]. Phosphorescence [Pri42]. Phosphors [Pri42]. Photo [GH45]. Photo-Disintegration [GH45]. Photochemical [Rab42]. Photographic [LR49, Sha41]. Photolysis [GI48]. Photons [Cla49a, Ros49b]. Photosynthetic [Fra45]. Physical [Bir41, Bob42, Kit49]. Physicist [Bee45, Eps48]. Physics [Cha43, GG46, Loo40, Mar41, Swa41a]. Planck [Lon45]. Plane [Gou45]. Planetary [Cle47]. Planets [Bob44, Wil42a]. Plastic [Bri45a]. Polar [PW48]. Polyatomic [Den40, Her42b, Sam46, Sha44, ST41]. Polymethine [HS42]. Polynomials [BW47]. Porphyrins [Rab44]. Postulate [Rob49]. Postulational [Bha49]. Potassium [KM42, Pri42]. Potassium-Halide [Pri42]. Potential [Bar49]. Power [SY49, Ske46]. Presence [Her42a]. Present [Mil49b]. Pressure [Bri45a, DP40, DP41]. Pressures [Bri46a, Bri46b, DG41]. Primary [BFL^{+}49, Meg42b]. Principle [Cas45]. Principles [Con42]. Problem [Wan45]. Problems [Cha43, Coc49, Loo40, Smy48b, Wil45, Smy48a]. Process [Kau42]. Processes [Mei45]. Produced [Kor49]. Production [Hei49, Kor48, Ske46, TF48]. Products [Ham48]. Progress [Str44].

Properties [BW45, Bri45a, Ham48, PE49]. Proportional [Kor42]. Proton [CL48]. Purposes [Smy45]. Pyridine [Spo44]. Pyroelectricity [Bor45].

Quantization [BB49, Pau43]. Quantum [Bec45, Bor45, Dir45, Fey48, KL45, PV49, Swa41b].

R. [DuB48]. Radial [Ske46]. Radiation [BFL^{+}49, Cla49a, FGV49, FLL48, Kor48, Kor49, Lat47, Val49, WF45]. Radio [Con42]. Radioactive [MBH47]. Radioactivities [LS40]. Radiofrequency [KM46]. Raman [Glo43]. Random [Gou45]. Rare [Meg42a]. Rate [Kau42, dHvdB49]. Ratio [PE48]. Ray [BMN^{+}48a, BMN^{+}48b, EE48, HW45, Hir42, KSW49, NR48, Opp49, Roc49, RG41, Ros48]. Rays [AAL^{+}48, AAC49, Gin43, Hev45, Lat47, MBH47, Ros49b]. Reaction [Dus45, TW49a]. Reactions [HTW49]. Reciprocity [Bor49]. Red [AB44, Bar42, Den40, GK45, NTD44, NN44, ORA44, OR44]. Reflection [Wil42c]. Refraction [Hel42]. Region [Ger49]. Regularization [PV49]. Related [MB40]. Relating [Bri45b]. Relation [Swa41a, Wil45]. Relations [Bar48, Cra48]. Relationships [The46, Ano47]. Relative [Bro49b]. Relativistic [Bha45, Dir49, Fey48, Gam49, PE49, Pau41, PV49, Tho45]. Relativity [Cle47, IS49, Kar44, Lem49, Rob49, Swa41b]. Relaxation [Kau42, PW48]. Release [Mei45]. Remarks [ES46, FL49, KL45, Opp49]. Report [GG46]. Research [Jew48, Wen47]. Resonance [Bro42, Fis42, KM46]. Rest [Ros49a]. Results [Coc49, Str49]. Reversibility [Cas45]. Reversible [GI48]. Review [PS48b]. Riemannian [Hof49, Lan49]. Rigid [Fok49]. Robert [Eps48]. Roentgen [Hev45]. Rollin [dHvdB49]. Rooms [MB44a, MB44b]. Rotation [Val49]. Rotational [TF42]. Rough [EES49]. Roughness [Bik44a, Bik44b]. Rubidium [KM42]. Rule [Win44].

Salts [MM42]. Satellites [Hir42]. Scalar [PH45]. Scalar- [PH45]. Scattering [Bar49]. Schemes [Mit48]. Science [Fra49]. Scientists [Fra41]. Search [Cra48]. Sections [GIF47]. Selection [Win44]. Self [CD48, EE48, PE49]. Self-Absorption [CD48, EE48]. Self-Energies [PE49]. Semi [Fee47]. Semi-Empirical [Fee47]. Sensitized [Win44]. Separation [JF46]. Series [CPS42]. Seventieth [Mil49a]. Shape [VW45]. Shell [GK45]. Shifts [Bar49]. Shock [BT49]. Showers [Cla49b, Coc49, MFDA49, Roc49]. Similiarity [vKL49]. Simon [dHvdB49]. Single [Mei45]. Sky [Elv42]. Sliding [Bik44a, Bik44b]. Slits [PB42]. Slowing [Mar47]. Small [BW45]. Solar [Cha46, FGV49, NR48]. Solid [KPC45, MB40]. Solution [Eck48a]. Solutions [Fre42, Göd49, KPC45]. Solving [Tau49]. Some [KL45, Mei45, Str49, Whe49]. Sorbed [GI48]. Sound [MB44a, MB44b]. Source [GK45]. Sources [EE48, Fis42, Liv46]. Space [ES45, ES46, Fey48, Fok49, Wil45]. Space-Time [Fey48, Wil45]. Spaces [Lan49]. Spark [Loe48]. Special [Tau49, Kar44, Rob49]. Species [Bro49b]. Specific [Lei48]. Spectra [Bar42, Bob44, CPS42, Den40, Fre42, Hal47, KM46, KM42, MB42, McK42, Meg42a, MR42, Mul42a, Mul42c, ORA44, OR44, PB42, Rab42, Rab44, Sam46, She42a, She42b, Skl42, ST41, Spo42, Str44, Swi42]. Spectroscopic [Bob42]. Spectroscopy [Boy41, Gor48, Lat47, Mei42, PK44]. Spectrum [AB44, Cla49a, CD48, Deu48, KSW49, Lei48, MM42, Mro42, NN44, TW49b]. Spin [PH45]. Spin-Dependent [PH45]. Square [BW47]. Standard [Meg42b]. Stars [ES45, GK45, ES46]. State [KL45]. Statements [Bri45b]. States [FLL48, NW49]. Static [Ros49a]. Statistical [Loe48, Wan45, Wea48]. Status [Mil49b]. Steel [Zen45]. Stellar [Cha44, Cha49a, Str44]. Stochastic [Cha43]. Stopped [SY49]. Storms [NR48]. Stress [Zen45]. Stresses [BW45]. Strong [PH45]. Structural [Van42]. Structure [Fis42, Mul42c, TF42]. Structures [Mul42a]. Studies [BMN^{+}48b, EE48]. Study [Cha49b]. Studying [MB42]. Sublimation [MR42]. Substances [PW48]. Substituted [Spo42]. Sudden [FGV49]. Summary [Hir42]. Sun [Bow48, Deu48]. Superconductive [KL45]. Superfluid [Dar40]. Supersonic [LB49]. Surface [Bee48, Bik44a, Bik44b, Fee47, Hen49]. Surfaces [Her49]. Surrounding [ES45, ES46]. Survey [Van45]. Symmetry [Fre42]. Symposium [Opp49]. System [Cha46]. Systems [Hel42, NW49].

Table [Bir41, Bro49b, LS40, Sea44, SP48]. Teacher [Eps48]. Teaching [Rol48]. Technique [Est46]. Technology [DuB48]. Telescope [BMN^{+}48a]. Television [Mor46]. tem [vL49]. Temperature [Lon45]. Tensors [Rac49]. Term [Ste49b]. Terms [WF49]. Terrestrial [Wil42a]. Test [IS49]. Texture [Van42]. Thallium [Pri42]. Thallium-Activated [Pri42]. Their [Mei45, Mul42c, Rab42]. Theories [Bar48, BB49, Pau41, PH45]. Theory [Bec45, Bha49, Bob42, Bor45, Bor49, Cha46, Dus45, Eck48b, Ein48, Fee47, Fra41, HW45, Hei49, Hul45, Kar44, Kit49, KL45, Mar47, PV49, Rob49, RG41, Swa41b, Swa41a, Van45, WU45, Wei49, Wen47, Wil45, Yuk49, vKL49, Str49]. Therefrom [DC48, DC49]. Thermal [JF46]. Thermionic [HN49]. Thermodynamical [Kle49]. Thin [Cha49b]. Time [Fey48, Fok49, PW48, Wil45]. Tracks [Sha41]. Transfer [Lon45, Rab42, dHvdB49]. Transformations [Sch49]. Triatomic [Bar42, Mul42a]. Tubes [Ske46]. Turbulence [vKL49]. Turbulent [LB49]. Type [Göd49, Her42b]. Types [Zwi40].

Ultraviolet [Boy41, Mul42c, PK44, Spo42, Spo44, TF42]. Undergraduate [Rol48]. Underwater [AY48]. Unified [Str49]. Universe [Ano49a]. Upper [Whi42, Whi43]. Uranium [Mei45]. Use [Hal47, Mit48, Wil42c]. Using [Mei45].

Vacuum [Boy41, PK44]. Valence [PE48]. Values [Bir41]. Vapor [DG41, OR44, Spo44]. Vapors [DG41]. Variable [BR45]. Various [Fre42]. Vector [PH45]. Vector-Pair [PH45]. Velocities [Dus45]. versus [Rob49]. Very [LR49]. Vibration [Sha44]. Vibrational [Hal47, Lei48]. Vibrations [Hou48]. Virtues [Bro45]. Voltage [Ske46].

Wall [EES49]. War [Jew48]. Wave [Bha45, Con42, Eck48a, Her49, Meg42b]. Wave-Length [Meg42b]. Waves [BT49, MB44a, MB44b]. Weaknesses [Bro45]. Weizsäcker [Cha46]. Width [AB44]. Widths [CL48]. Will [Wea48]. Wilson [GG46]. Work [Bri46a, Bri46b, Jew48]. World [Jew48]. Wurster [MM42].

X [Gin43, Hir42]. X-Ray [Hir42]. X-Rays [Gin43].

Yields [Ano47, The46].

Zone [Hen49].

References

Adams:1949:OCR

[AAC49] Raymond V. Adams, Carl D. Anderson, and Eugene W. Cowan. Observations of cosmic rays at high altitudes. Reviews of Modern Physics, 21(1):72–75, January 1949. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

           0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
           21.72;   http://rmp.aps.org/abstract/RMP/v21/i1/p72_1.

Adams:1948:CRF

[AAL^{+}48] Raymond V. Adams, Carl D. Anderson, Paul E. Lloyd, R. Ronald Rau, and Ram C. Saxena. Cosmic rays at 30,000 feet. Reviews of Modern Physics, 20(1):334–349, January 1948. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

           URL http://link.aps.org/doi/10.1103/RevModPhys.20.334;
            http://rmp.aps.org/abstract/RMP/v20/i1/p334_1.

Adel:1944:GIR

[AB44] Arthur Adel and E. F. Barker. Grating infra-red measurements at oblique incidence. Line width in the spectrum of N_{2}O. Reviews of Modern Physics, 16(3–4):236–240, July 1944. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

           URL http://link.aps.org/doi/10.1103/RevModPhys.16.236;
            http://rmp.aps.org/abstract/RMP/v16/i3-4/p236_1.

Anonymous:1942:Da

[Ano42a] Anonymous. Discussion. Reviews of Modern Physics, 14 (2–3):63, April 1942. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.14.63;   http://
rmp.aps.org/abstract/RMP/v14/i2-3/p63_1.

Anonymous:1942:Db

[Ano42b] Anonymous. Discussion. Reviews of Modern Physics, 14 (2–3):65, April 1942. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.14.65;   http://
rmp.aps.org/abstract/RMP/v14/i2-3/p65_1.

Anonymous:1942:Dc

[Ano42c] Anonymous. Discussion. Reviews of Modern Physics, 14(2–3):150, April 1942. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.14.150;   http://rmp.aps.org/abstract/
RMP/v14/i2-3/p150_1.

Anonymous:1942:Dd

[Ano42d] Anonymous. Discussion. Reviews of Modern Physics, 14(2–3):178, April 1942. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.14.178;   http://rmp.aps.org/abstract/
RMP/v14/i2-3/p178_1.

Anonymous:1942:De

[Ano42e] Anonymous. Discussion. Reviews of Modern Physics, 14(2–3):189, April 1942. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.14.189;   http://rmp.aps.org/abstract/
RMP/v14/i2-3/p189_1.

Anonymous:1942:Df

[Ano42f] Anonymous. Discussion. Reviews of Modern Physics, 14(2– 3):257–258, April 1942. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.14.257;   http://
rmp.aps.org/abstract/RMP/v14/i2-3/p257_1.

Anonymous:1947:ENF

[Ano47] Anonymous. Erratum: Nuclei Formed in Fission: Decay Characteristics, Fission Yields, and Chain Relationships. Reviews of Modern Physics, 19(1):18, January 1947. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.19.18;   http://
rmp.aps.org/abstract/RMP/v19/i1/p18_1. See [The46].

Anonymous:1949:AU

[Ano49a] Anonymous. The age of the universe. Reviews of Modern Physics, 21(3):374–378, July 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.21.374;   http:/
/rmp.aps.org/abstract/RMP/v21/i3/p374_1. Special issue in

honor of Albert Einstein on his 70th birthday.

Anonymous:1949:D

[Ano49b] Anonymous. Discussion. Reviews of Modern Physics, 21 (1):49–50, January 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.21.49;   http://
rmp.aps.org/abstract/RMP/v21/i1/p49_1.

Arons:1948:EPU

[AY48] A. B. Arons and D. R. Yennie. Energy partition in underwater explosion phenomena. Reviews of Modern Physics, 20(3):519–536, July 1948. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.20.519;   http://rmp.aps.org/abstract/
RMP/v20/i3/p519_1.

Barker:1942:IRS

[Bar42] Ernest F. Barker. The infra-red spectra of triatomic molecules. Reviews of Modern Physics, 14(2–3):198–203, April 1942. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
14.198;   http://rmp.aps.org/abstract/RMP/v14/i2-3/p198_
1.

Barnett:1948:NDC

[Bar48] S. J. Barnett. Note on diamagnetism: Clarification of relations between fundamental molecular theories. Reviews of Modern Physics, 20(1):78–81, January 1948. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
20.78;   http://rmp.aps.org/abstract/RMP/v20/i1/p78_1.

Bargmann:1949:CBP

[Bar49] V. Bargmann. On the connection between phase shifts and scattering potential. Reviews of Modern Physics, 21(3):488–493, July 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.21.488;   http://rmp.aps.org/abstract/

RMP/v21/i3/p488_1. Special issue in honor of Albert Einstein on his 70th birthday.

Bergmann:1949:NLF

[BB49] Peter G. Bergmann and Johanna H. M. Brunings. Non-linear field theories II. Canonical equations and quantization. Reviews of Modern Physics, 21(3):480–487, July 1949. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.21.480;
 http://rmp.aps.org/abstract/RMP/v21/i3/p480_1. Special

issue in honor of Albert Einstein on his 70th birthday.

Bartky:1948:AEF

[BD48] W. Bartky and A. J. Dempster. The approach to equilibrium in fractionation. Reviews of Modern Physics, 20(1):123–127, January 1948. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.20.123;   http://rmp.aps.org/abstract/
RMP/v20/i1/p123_1.

Beck:1945:FCQ

[Bec45] Guido Beck. Field concepts in quantum theory. Reviews of Modern Physics, 17(2–3):187–194, April 1945. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.17.187;
 http://rmp.aps.org/abstract/RMP/v17/i2-3/p187_1.

Beeck:1945:CCP

[Bee45] Otto Beeck. Catalysis — a challenge to the physicist (as exemplified by the hydrogenation of ethylene over metal catalysts). Reviews of Modern Physics, 17(1):61–71, January 1945. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
17.61;   http://rmp.aps.org/abstract/RMP/v17/i1/p61_1.

Beeck:1948:SC

[Bee48] Otto Beeck. Surface catalysis. Reviews of Modern Physics, 20 (1):127–130, January 1948. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.20.127;   http://
rmp.aps.org/abstract/RMP/v20/i1/p127_1.

Bradt:1949:EHN

[BFL^{+}49] H. L. Bradt, Phyllis Freier, E. J. Lofgren, E. P. Ney, F. Oppenheimer, and B. Peters. Evidence for heavy nuclei as a component of primary cosmic radiation. Reviews of Modern Physics, 21(1):101–103, January 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.21.101;   http://
rmp.aps.org/abstract/RMP/v21/i1/p101_1.

Bhabha:1945:RWE

[Bha45] H. J. Bhabha. Relativistic wave equations for the elementary particles. Reviews of Modern Physics, 17(2–3):200–216, April 1945. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.17.200;   http://rmp.aps.org/abstract/
RMP/v17/i2-3/p200_1.

Bhabha:1949:PBT

[Bha49] H. J. Bhabha. On the postulational basis of the theory of elementary particles. Reviews of Modern Physics, 21(3):451–462, July 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.21.451;   http://rmp.aps.org/abstract/

RMP/v21/i3/p451_1. Special issue in honor of Albert Einstein on his 70th birthday.

Bikerman:1944:SRSa

[Bik44a] J. J. Bikerman. Surface roughness and sliding friction. Reviews of Modern Physics, 16(1):53–68, January 1944. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
16.53;   http://rmp.aps.org/abstract/RMP/v16/i1/p53_1.

Bikerman:1944:SRSb

[Bik44b] J. J. Bikerman. Surface roughness and sliding friction. Reviews of Modern Physics, 16(3–4):324, July 1944. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.16.324;
 http://rmp.aps.org/abstract/RMP/v16/i3-4/p324_2.

Birge:1941:NTV

[Bir41] Raymond T. Birge. A new table of values of the general physical constants (as of August, 1941). Reviews of Modern Physics, 13(4):233–239, October 1941. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.13.233;   http://
rmp.aps.org/abstract/RMP/v13/i4/p233_1.

Biehl:1948:NCR

[BMN^{+}48a] A. T. Biehl, R. A. Montgomery, H. V. Neher, W. H. Pickering, and W. C. Roesch. A new cosmic-ray telescope for high altitudes. Reviews of Modern Physics, 20(1):353–359, January 1948. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
20.353;   http://rmp.aps.org/abstract/RMP/v20/i1/p353_
1.

Biehl:1948:RSC

[BMN^{+}48b] A. T. Biehl, R. A. Montgomery, H. V. Neher, W. H. Pickering, and W. C. Roesch. Recent studies of the cosmic-ray latitude effect at high altitudes. Reviews of Modern Physics, 20(1):360–366, January 1948. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.20.360;   http://rmp.aps.org/abstract/
RMP/v20/i1/p360_1.

Bobrovnikoff:1942:PTC

[Bob42] N. T. Bobrovnikoff. Physical theory of comets in the light of spectroscopic data. Reviews of Modern Physics, 14(2–3):164–178, April 1942. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.14.164;   http://rmp.aps.org/abstract/
RMP/v14/i2-3/p164_1.

Bobrovnikoff:1944:SP

[Bob44] N. T. Bobrovnikoff. Spectra of the planets. Reviews of Modern Physics, 16(3–4):271–285, July 1944. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.16.271;
 http://rmp.aps.org/abstract/RMP/v16/i3-4/p271_1.

Born:1945:QTP

[Bor45] Max Born. On the quantum theory of pyroelectricity. Reviews of Modern Physics, 17(2–3):245–251, April 1945. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.17.245;
 http://rmp.aps.org/abstract/RMP/v17/i2-3/p245_1.

Born:1949:RTE

[Bor49] Max Born. Reciprocity theory of elementary particles. Reviews of Modern Physics, 21(3):463–473, July 1949. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.21.463;
 http://rmp.aps.org/abstract/RMP/v21/i3/p463_1. Special

issue in honor of Albert Einstein on his 70th birthday.

Bowen:1948:AOS

[Bow48] I. S. Bowen. The abundance of oxygen in the sun. Reviews of Modern Physics, 20(1):109–112, January 1948. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.20.109;
 http://rmp.aps.org/abstract/RMP/v20/i1/p109_1.

Boyce:1941:SVU

[Boy41] J. C. Boyce. Spectroscopy in the vacuum ultraviolet. Reviews of Modern Physics, 13(1):1–57, January 1941. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.13.1;
http://rmp.aps.org/abstract/RMP/v13/i1/p1_1.

Bozorth:1947:M

[Boz47] R. M. Bozorth. Magnetism. Reviews of Modern Physics, 19 (1):29–86, January 1947. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.19.29;   http://
rmp.aps.org/abstract/RMP/v19/i1/p29_1.

Bloch:1945:AVM

[BR45] F. Bloch and I. I. Rabi. Atoms in variable magnetic fields. Reviews of Modern Physics, 17(2–3):237–244, April 1945. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.17.237;
 http://rmp.aps.org/abstract/RMP/v17/i2-3/p237_1.

Bridgman:1945:EHH

[Bri45a] P. W. Bridgman. Effects of high hydrostatic pressure on the plastic properties of metals. Reviews of Modern Physics, 17 (1):3–14, January 1945. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.17.3;   http://rmp.aps.
org/abstract/RMP/v17/i1/p3_1.

BSI:1945:SRA

[Bri45b] British Information Services. Statements relating to the atomic bomb. Reviews of Modern Physics, 17(4):472–490, October 1945. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.17.472;   http://rmp.aps.org/abstract/
RMP/v17/i4/p472_1.

Bridgman:1946:RWFa

[Bri46a] P. W. Bridgman. Recent work in the field of high pressures. Reviews of Modern Physics, 18(1):1–93, January 1946. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.18.1;
http://rmp.aps.org/abstract/RMP/v18/i1/p1_1. See adden-

dum [Bri46b].

Bridgman:1946:RWFb

[Bri46b] P. W. Bridgman. Recent work in the field of high pressures: Addendum. Reviews of Modern Physics, 18(2):291, April 1946. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
18.291;   http://rmp.aps.org/abstract/RMP/v18/i2/p291_

1. See [Bri46a].

Brooker:1942:ARD

[Bro42] Leslie G. S. Brooker. Absorption and resonance in dyes. Reviews of Modern Physics, 14(2–3):275–293, April 1942. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.14.275;
 http://rmp.aps.org/abstract/RMP/v14/i2-3/p275_1.

Brown:1945:VWD

[Bro45] William F. Brown, Jr. Virtues and weaknesses of the domain concept. Reviews of Modern Physics, 17(1):15–19, January 1945. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.17.15;   http://rmp.aps.org/abstract/RMP/v17/
i1/p15_1.

Brode:1949:MM

[Bro49a] Robert B. Brode. The mass of the mesotron. Reviews of Modern Physics, 21(1):37–41, January 1949. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
21.37;   http://rmp.aps.org/abstract/RMP/v21/i1/p37_1.

Brown:1949:TRA

[Bro49b] Harrison Brown. A table of relative abundances of nuclear species. Reviews of Modern Physics, 21(4):625–634, October 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
21.625;   http://rmp.aps.org/abstract/RMP/v21/i4/p625_
1.

Bleakney:1949:ISW

[BT49] Walker Bleakney and A. H. Taub. Interaction of shock waves. Reviews of Modern Physics, 21(4):584–605, October 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
21.584;   http://rmp.aps.org/abstract/RMP/v21/i4/p584_
1.

Bozorth:1945:ESS

[BW45] R. M. Bozorth and H. J. Williams. Effect of small stresses on magnetic properties. Reviews of Modern Physics, 17(1): 72–80, January 1945. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.17.72;   http://
rmp.aps.org/abstract/RMP/v17/i1/p72_1.

Birge:1947:LSF

[BW47] Raymond T. Birge and J. W. Weinberg. Least-squares’ fitting of data by means of polynomials. Reviews of Modern Physics, 19(4):298–360, October 1947. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.19.298;   http://
rmp.aps.org/abstract/RMP/v19/i4/p298_1.

Casimir:1945:OPM

[Cas45] H. B. G. Casimir. On Onsager’s principle of microscopic reversibility. Reviews of Modern Physics, 17(2–3):343–350, April 1945. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.17.343;   http://rmp.aps.org/abstract/
RMP/v17/i2-3/p343_1.

Cowan:1948:SAS

[CD48] R. D. Cowan and G. H. Dieke. Self-absorption of spectrum lines. Reviews of Modern Physics, 20(2):418–455, April 1948. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
20.418;   http://rmp.aps.org/abstract/RMP/v20/i2/p418_
1.

Chandrasekhar:1943:SPP

[Cha43] S. Chandrasekhar. Stochastic problems in physics and astronomy. Reviews of Modern Physics, 15(1):1–89, January 1943. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
15.1;   http://rmp.aps.org/abstract/RMP/v15/i1/p1_1.

Chandrasekhar:1944:NIH

[Cha44] S. Chandrasekhar. The negative ions of hydrogen and oxygen in stellar atmospheres. Reviews of Modern Physics, 16(3–4):301–306, July 1944. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.16.301;   http://rmp.aps.org/abstract/
RMP/v16/i3-4/p301_1.

Chandrasekhar:1945:FAL

[Cha45] S. Chandrasekhar. The formation of absorption lines in a moving atmosphere. Reviews of Modern Physics, 17(2–3):138–156, April 1945. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.17.138;   http://rmp.aps.org/abstract/
RMP/v17/i2-3/p138_1.

Chandrasekhar:1946:NTW

[Cha46] S. Chandrasekhar. On a new theory of Weizsäcker on the origin of the solar system. Reviews of Modern Physics, 18 (1):94–102, January 1946. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.18.94;   http://
rmp.aps.org/abstract/RMP/v18/i1/p94_1.

Chandrasekhar:1949:BMD

[Cha49a] S. Chandrasekhar. Brownian motion, dynamical friction, and stellar dynamics. Reviews of Modern Physics, 21(3):383–388, July 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.21.383;   http://rmp.aps.org/abstract/

RMP/v21/i3/p383_1. Special issue in honor of Albert Einstein on his 70th birthday.

Chang:1949:CCS

[Cha49b] W. Y. Chang. A cloud-chamber study of meson absorption by thin Pb, Fe, and Al foils. Reviews of Modern Physics, 21 (1):166–180, January 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.21.166;   http://
rmp.aps.org/abstract/RMP/v21/i1/p166_1.

Christy:1948:PWL

[CL48] R. F. Christy and R. Latter. Proton widths of light nuclei. Reviews of Modern Physics, 20(1):185–190, January 1948. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.20.185;
 http://rmp.aps.org/abstract/RMP/v20/i1/p185_1.

Clay:1949:ESC

[Cla49a] J. Clay. The energy spectrum of cosmic-radiation particles and photons. Reviews of Modern Physics, 21(1):82–93, January 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.21.82;   http://rmp.aps.org/abstract/RMP/v21/
i1/p82_1.

Clay:1949:HEP

[Cla49b] J. Clay. High energy particles, bursts, and showers. Reviews of Modern Physics, 21(1):94–100, January 1949. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
21.94;   http://rmp.aps.org/abstract/RMP/v21/i1/p94_1.

Clemence:1947:REP

[Cle47] G. M. Clemence. The relativity effect in planetary motions. Reviews of Modern Physics, 19(4):361–364, October 1947. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
19.361;   http://rmp.aps.org/abstract/RMP/v19/i4/p361_
1.

Cocconi:1949:RPC

[Coc49] G. Cocconi. Results and problems concerning the extensive air showers. Reviews of Modern Physics, 21(1):26–30, January 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.21.26;   http://rmp.aps.org/abstract/RMP/v21/
i1/p26_1.

Condon:1942:PMW

[Con42] E. U. Condon. Principles of micro-wave radio. Reviews of Modern Physics, 14(4):341–389, October 1942. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.14.341;
 http://rmp.aps.org/abstract/RMP/v14/i4/p341_1.

Constant:1945:FIM

[Con45] F. W. Constant. Ferromagnetic impurities in metals. Reviews of Modern Physics, 17(1):81–86, January 1945. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
17.81;   http://rmp.aps.org/abstract/RMP/v17/i1/p81_1.

Carr:1942:ASS

[CPS42] Emma P. Carr, Lucy W. Pickett, and Hildegard Stücklen. The absorption spectra of a series of dienes. Reviews of Modern Physics, 14(2–3):260–264, April 1942. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.14.260;
 http://rmp.aps.org/abstract/RMP/v14/i2-3/p260_1.

Crane:1948:EMR

[Cra48] H. R. Crane. The energy and momentum relations in the betadecay, and the search for the neutrino. Reviews of Modern Physics, 20(1):278–295, January 1948. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.20.278;   http://
rmp.aps.org/abstract/RMP/v20/i1/p278_1.

Darrow:1940:HS

[Dar40] Karl K. Darrow. Helium the superfluid. Reviews of Modern Physics, 12(3):257–266, July 1940. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.12.257;   http:/
/rmp.aps.org/abstract/RMP/v12/i3/p257_1.

Darrow:1944:F

[Dar44] Karl K. Darrow. Foreword. Reviews of Modern Physics, 16(3–4):152, July 1944. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.16.152;   http://
rmp.aps.org/abstract/RMP/v16/i3-4/p152_1.

Darrow:1945:F

[Dar45] Karl K. Darrow. Foreword. Reviews of Modern Physics, 17 (1):2, January 1945. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.17.2;   http://rmp.aps.
org/abstract/RMP/v17/i1/p2_1.

deBroglie:1949:LED

[dB49] L. de Broglie. L’oeuvre d’Einstein et la dualité des ondes et des corpuscules. Reviews of Modern Physics, 21(3):345–347, July 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.21.345;   http://rmp.aps.org/abstract/

RMP/v21/i3/p345_1. Special issue in honor of Albert Einstein on his 70th birthday.

DuMond:1948:OKA

[DC48] Jesse W. M. DuMond and E. Richard Cohen. Our knowledge of the atomic constants F, N, m, and h in 1947, and of other constants derivable therefrom. Reviews of Modern Physics, 20 (1):82–108, January 1948. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.20.82;   http://
rmp.aps.org/abstract/RMP/v20/i1/p82_1. See [DC49].

DuMond:1949:OKA

[DC49] J. W. M. DuMond and E. Richard Cohen. Our knowledge of the atomic constants F, N, m, and h in 1947, and of other constants derivable therefrom. Reviews of Modern Physics, 21 (4):651–652, October 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.21.651;   http://
rmp.aps.org/abstract/RMP/v21/i4/p651_1. See [DC48].

Dennison:1940:IRS

[Den40] David M. Dennison. The infra-red spectra of polyatomic molecules. Part II. Reviews of Modern Physics, 12(3):175–214, July 1940. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.12.175;   http://rmp.aps.org/abstract/
RMP/v12/i3/p175_1.

Deutsch:1948:CSS

[Deu48] Armin J. Deutsch. The continuous spectrum of the sun. Reviews of Modern Physics, 20(2):388–398, April 1948. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.20.388;
 http://rmp.aps.org/abstract/RMP/v20/i2/p388_1.

Ditchburn:1941:VPM

[DG41] R. W. Ditchburn and J. C. Gilmour. The vapor pressures of monatomic vapors. Reviews of Modern Physics, 13(4):310–327, October 1941. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.13.310;   http://rmp.aps.org/abstract/
RMP/v13/i4/p310_1.

deHaas:1949:TRR

[dHvdB49] W. J. de Haas and G. J. van den Berg. On the transfer rate of the Rollin–Simon film. Reviews of Modern Physics, 21(3):524–526, July 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.21.524;   http://rmp.aps.org/abstract/

RMP/v21/i3/p524_1. Special issue in honor of Albert Einstein on his 70th birthday.

Dirac:1945:ABC

[Dir45] P. A. M. Dirac. On the analogy between classical and quantum mechanics. Reviews of Modern Physics, 17(2–3):195–199, April 1945. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.17.195;   http://rmp.aps.org/abstract/
RMP/v17/i2-3/p195_1.

Dirac:1949:FRD

[Dir49] P. A. M. Dirac. Forms of relativistic dynamics. Reviews of Modern Physics, 21(3):392–399, July 1949. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.21.392;
 http://rmp.aps.org/abstract/RMP/v21/i3/p392_1. Special

issue in honor of Albert Einstein on his 70th birthday.

Druyvesteyn:1940:MED

[DP40] M. J. Druyvesteyn and F. M. Penning. The mechanism of electrical discharges in gases of low pressure. Reviews of Modern Physics, 12(2):87–174, April 1940. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.12.87;   http://
rmp.aps.org/abstract/RMP/v12/i2/p87_1. See errata [DP41].

Druyvesteyn:1941:EME

[DP41] M. J. Druyvesteyn and F. M. Penning. Errata: Mechanism of Electrical Discharges in Gases of Low Pressure. Reviews of Modern Physics, 13(1):72–73, January 1941. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.13.72;   http://
rmp.aps.org/abstract/RMP/v13/i1/p72_1. See [DP40].

DuBridge:1948:RMC

[DuB48] L. A. DuBridge. R. A. Millikan and the California Institute of Technology. Reviews of Modern Physics, 20(1):7–9, January 1948. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.

org/doi/10.1103/RevModPhys.20.7;   http://rmp.aps.org/
abstract/RMP/v20/i1/p7_1.

Dushman:1945:ATA

[Dus45] Saul Dushman. Application of theory of absolute reaction velocities to creep of metals. Reviews of Modern Physics, 17 (1):48–49, January 1945. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.17.48;   http://
rmp.aps.org/abstract/RMP/v17/i1/p48_1.

Eckart:1948:ASO

[Eck48a] Carl Eckart. The approximate solution of one-dimensional wave equations. Reviews of Modern Physics, 20(2):399–417, April 1948. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.20.399;   http://rmp.aps.org/abstract/
RMP/v20/i2/p399_1.

Eckart:1948:TAF

[Eck48b] Carl Eckart. The theory of the anelastic fluid. Reviews of Modern Physics, 20(1):232–235, January 1948. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.20.232;
 http://rmp.aps.org/abstract/RMP/v20/i1/p232_1.

Evans:1948:SSA

[EE48] Robley D. Evans and Richard O. Evans. Studies of selfabsorption in gamma-ray sources. Reviews of Modern Physics, 20(1):305–326, January 1948. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.20.305;   http://
rmp.aps.org/abstract/RMP/v20/i1/p305_1.

Einstein:1949:HFR

[EES49] Hans Albert Einstein and El-Sayed Ahmed El-Samni. Hydrodynamic forces on a rough wall. Reviews of Modern Physics, 21(3):520–524, July 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.21.520;   http://
rmp.aps.org/abstract/RMP/v21/i3/p520_1. Special issue in

honor of Albert Einstein on his 70th birthday.

Einstein:1948:GTG

[Ein48] Albert Einstein. A generalized theory of gravitation. Reviews of Modern Physics, 20(1):35–39, January 1948. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/abstract/RMP/v20/p35;   http:
//link.aps.org/doi/10.1103/RevModPhys.20.35;   http://
prola.aps.org/pdf/RMP/v20/i1/p35_1;   http://rmp.aps.org/
abstract/RMP/v20/i1/p35_1.

Eliezer:1947:IEE

[Eli47] C. Jayaratnam Eliezer. The interaction of electrons and an electromagnetic field. Reviews of Modern Physics, 19(3):147–184, July 1947. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.19.147;   http://rmp.aps.org/abstract/
RMP/v19/i3/p147_1.

Elvey:1942:LNS

[Elv42] C. T. Elvey. The light of the night sky. Reviews of Modern Physics, 14(2–3):140–150, April 1942. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.14.140;   http:/
/rmp.aps.org/abstract/RMP/v14/i2-3/p140_1.

Epstein:1948:RAM

[Eps48] Paul S. Epstein. Robert Andrews Millikan as physicist and teacher. Reviews of Modern Physics, 20(1):10–25, January 1948. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
20.10;   http://rmp.aps.org/abstract/RMP/v20/i1/p10_1.

Einstein:1945:IES

[ES45] Albert Einstein and Ernst G. Straus. The influence of the expansion of space on the gravitation fields surrounding the individual stars. Reviews of Modern Physics, 17(2– 3):120–124, April 1945. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/abstract/RMP/v17/p120;   http://link.aps.
org/doi/10.1103/RevModPhys.17.120;   http://prola.aps.
org/pdf/RMP/v17/i2-3/p120_1;   http://rmp.aps.org/abstract/

RMP/v17/i2-3/p120_1. See corrections and additional remarks [ES46].

Einstein:1946:CAR

[ES46] A. Einstein and E. G. Straus. Corrections and additional remarks to our paper: The Influence of the Expansion of Space on the Gravitation Fields Surrounding the Individual Stars. Reviews of Modern Physics, 18(1):148–149, January 1946. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/abstract/RMP/v18/p148;   http:
//link.aps.org/doi/10.1103/RevModPhys.18.148;   http:/
/rmp.aps.org/abstract/RMP/v18/i1/p148_1. See [ES45].

Estermann:1946:MBT

[Est46] I. Estermann. Molecular beam technique. Reviews of Modern Physics, 18(3):300–323, July 1946. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.18.300;   http:/
/rmp.aps.org/abstract/RMP/v18/i3/p300_1.

Feenberg:1947:SET

[Fee47] Eugene Feenberg. Semi-empirical theory of the nuclear energy surface. Reviews of Modern Physics, 19(3):239–258, July 1947. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.19.239;   http://rmp.aps.org/abstract/
RMP/v19/i3/p239_1.

Feynman:1948:STA

[Fey48] R. P. Feynman. Space-time approach to non-relativistic quantum mechanics. Reviews of Modern Physics, 20(2):367–387, April 1948. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.20.367;   http://rmp.aps.org/abstract/
RMP/v20/i2/p367_1.

Forbush:1949:MSI

[FGV49] S. E. Forbush, P. S. Gill, and M. S. Vallarta. On the mechanism of sudden increases of cosmic radiation associated with solar flares. Reviews of Modern Physics, 21(1):44–48, January 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
21.44;   http://rmp.aps.org/abstract/RMP/v21/i1/p44_1.

Field:1946:HCE

[Fie46] L. M. Field. High current electron guns. Reviews of Modern Physics, 18(3):353–361, July 1946. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.18.353;   http:/
/rmp.aps.org/abstract/RMP/v18/i3/p353_1.

Fisher:1942:ABL

[Fis42] Russell A. Fisher. Atomic beam light sources applied to the structure of the magnesium I resonance line. Reviews of Modern Physics, 14(2–3):79–81, April 1942. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.14.79;   http://
rmp.aps.org/abstract/RMP/v14/i2-3/p79_1.

Franck:1949:RII

[FL49] James Franck and Robert Livingston. Remarks on intraand inter-molecular migration of excitation energy. Reviews of Modern Physics, 21(3):505–509, July 1949. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.21.505;
 http://rmp.aps.org/abstract/RMP/v21/i3/p505_1. Special

issue in honor of Albert Einstein on his 70th birthday.

Fletcher:1940:AP

[Fle40] Harvey Fletcher. Auditory patterns. Reviews of Modern Physics, 12(1):47–65, January 1940. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.12.47;   http://
rmp.aps.org/abstract/RMP/v12/i1/p47_1.

Fowler:1948:GRE

[FLL48] W. A. Fowler, C. C. Lauritsen, and T. Lauritsen. Gamma-radiation from excited states of light nuclei. Reviews of Modern Physics, 20(1):236–277, January 1948. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.20.236;   http://
rmp.aps.org/abstract/RMP/v20/i1/p236_1.

Fokker:1949:STG

[Fok49] A. D. Fokker. On the space–time geometry of a moving rigid body. Reviews of Modern Physics, 21(3):406–408, July 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
21.406;   http://rmp.aps.org/abstract/RMP/v21/i3/p406_

1. Special issue in honor of Albert Einstein on his 70th birthday.

Frank:1941:WDS

[Fra41] Philipp Frank. Why do scientists and philosophers so often disagree about the merits of a new theory? Reviews of Modern Physics, 13(3):171–175, July 1941. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.13.171;
 http://rmp.aps.org/abstract/RMP/v13/i3/p171_1.

Franck:1945:PAI

[Fra45] James Franck. Photosynthetic activity of isolated chloroplasts. Reviews of Modern Physics, 17(2–3):112–119, April 1945. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.17.112;   http://rmp.aps.org/abstract/
RMP/v17/i2-3/p112_1.

Frank:1949:EPS

[Fra49] Philipp Frank. Einstein’s philosophy of science. Reviews of Modern Physics, 21(3):349–355, July 1949. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.21.349;
 http://rmp.aps.org/abstract/RMP/v21/i3/p349_1. Special

issue in honor of Albert Einstein on his 70th birthday.

Freed:1942:SIF

[Fre42] Simon Freed. Spectra of ions in fields of various symmetry in crystals and solutions. Reviews of Modern Physics, 14(2–3):105–111, April 1942. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.14.105;   http://rmp.aps.org/abstract/
RMP/v14/i2-3/p105_1.

Gamow:1949:RC

[Gam49] G. Gamow. On relativistic cosmogony. Reviews of Modern Physics, 21(3):367–373, July 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.21.367;   http:/
/rmp.aps.org/abstract/RMP/v21/i3/p367_1. Special issue in

honor of Albert Einstein on his 70th birthday.

Gerson:1949:NIF

[Ger49] N. C. Gerson. Nocturnal ionization in the F_{2} ionospheric region. Reviews of Modern Physics, 21(4):606–624, October 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
21.606;   http://rmp.aps.org/abstract/RMP/v21/i4/p606_
1.

Gupta:1946:RWC

[GG46] N. N. Das Gupta and S. K. Ghosh. A report on the Wilson cloud chamber and its applications in physics. Reviews of Modern Physics, 18(2):225–290, April 1946. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.18.225;
 http://rmp.aps.org/abstract/RMP/v18/i2/p225_1.

Graham:1945:ADN

[GH45] Gordon A. R. Graham and Hans Halban, Jr. On the angular distribution of neutrons in the photo-disintegration of the deuteron. Reviews of Modern Physics, 17(2–3):297–304, April 1945. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.17.297;   http://rmp.aps.org/abstract/
RMP/v17/i2-3/p297_1.

Goetz:1948:RPA

[GI48] Alexander Goetz and Edward C. Y. Inn. Reversible photolysis of Ag sorbed on collodial metal oxides. Reviews of Modern Physics, 20(1):131–142, January 1948. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.20.131;
 http://rmp.aps.org/abstract/RMP/v20/i1/p131_1.

Goldsmith:1947:NCS

[GIF47] H. H. Goldsmith, H. W. Ibser, and B. T. Feld. Neutron cross sections of the elements: a compilation. Reviews of Modern Physics, 19(4):259–297, October 1947. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.19.259;
 http://rmp.aps.org/abstract/RMP/v19/i4/p259_1.

Gingrich:1943:DXR

[Gin43] Newell S. Gingrich. The diffraction of X-rays by liquid elements. Reviews of Modern Physics, 15(1):90–110, January 1943. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
15.90;   http://rmp.aps.org/abstract/RMP/v15/i1/p90_1.

Gamow:1945:SSM

[GK45] G. Gamow and G. Keller. A shell source model for red giant stars. Reviews of Modern Physics, 17(2–3):125–137, April 1945. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.17.125;   http://rmp.aps.org/abstract/
RMP/v17/i2-3/p125_1.

Glockler:1943:RE

[Glo43] Geo. Glockler. The Raman effect. Reviews of Modern Physics, 15(2):111–173, April 1943. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.15.111;   http://
rmp.aps.org/abstract/RMP/v15/i2/p111_1.

Godel:1949:ENT

[Göd49] Kurt Gödel. An example of a new type of cosmological solutions of Einstein’s field equations of gravitation. Reviews of Modern Physics, 21(3):447–450, July 1949. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.21.447;
 http://rmp.aps.org/abstract/RMP/v21/i3/p447_1. Special

issue in honor of Albert Einstein on his 70th birthday.

Gordy:1948:MS

[Gor48] Walter Gordy. Microwave spectroscopy. Reviews of Modern Physics, 20(4):668–717, October 1948. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.20.668;   http:/
/rmp.aps.org/abstract/RMP/v20/i4/p668_1.

Goudsmit:1945:RDL

[Gou45] S. Goudsmit. Random distribution of lines in a plane. Reviews of Modern Physics, 17(2–3):321–322, April 1945. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.17.321;
 http://rmp.aps.org/abstract/RMP/v17/i2-3/p321_1.

Gutmann:1948:E

[Gut48] F. Gutmann. The electret. Reviews of Modern Physics, 20 (3):457–472, July 1948. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.20.457;   http://
rmp.aps.org/abstract/RMP/v20/i3/p457_1.

Halverson:1947:UDA

[Hal47] Frederick Halverson. The use of deuterium in the analysis of vibrational spectra. Reviews of Modern Physics, 19 (2):87–131, April 1947. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.19.87;   http://
rmp.aps.org/abstract/RMP/v19/i2/p87_1.

Hamilton:1948:MPF

[Ham48] Joseph G. Hamilton. The metabolic properties of the fission products and actinide elements. Reviews of Modern Physics, 20 (4):718–728, October 1948. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.20.718;   http://
rmp.aps.org/abstract/RMP/v20/i4/p718_1.

Hearmon:1946:ECA

[Hea46] R. F. S. Hearmon. The elastic constants of anisotropic materials. Reviews of Modern Physics, 18(3):409–440, July 1946. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
18.409;   http://rmp.aps.org/abstract/RMP/v18/i3/p409_
1.

Heitler:1949:TMP

[Hei49] W. Heitler. Theory of meson production. Reviews of Modern Physics, 21(1):113–121, January 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.21.113;   http:/
/rmp.aps.org/abstract/RMP/v21/i1/p113_1.

Heller:1942:OCE

[Hel42] Wilfried Heller. The origin and the complications of electric double refraction and of electric dichroism in dilute dispersed systems. Reviews of Modern Physics, 14(4):390–409, October 1942. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
14.390;   http://rmp.aps.org/abstract/RMP/v14/i4/p390_

Henniker:1949:DSZ

[Hen49] J. C. Henniker. The depth of the surface zone of a liquid. Reviews of Modern Physics, 21(2):322–341, April 1949. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.21.322;
 http://rmp.aps.org/abstract/RMP/v21/i2/p322_1.

Herzberg:1942:EPC

[Her42a] G. Herzberg. Evidence for the presence of CH_{2} molecules in comets. Reviews of Modern Physics, 14(2–3):195–197, April 1942. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.14.195;   http://rmp.aps.org/abstract/
RMP/v14/i2-3/p195_1.

Herzberg:1942:TDL

[Her42b] G. Herzberg. l-type doubling in linear polyatomic molecules. Reviews of Modern Physics, 14(2–3):219–223, April 1942. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
14.219;   http://rmp.aps.org/abstract/RMP/v14/i2-3/p219_
1.

Herzfeld:1949:NSM

[Her49] Karl F. Herzfeld. Nodal surfaces in molecular wave functions. Reviews of Modern Physics, 21(3):527–530, July 1949. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.21.527;
 http://rmp.aps.org/abstract/RMP/v21/i3/p527_1. Special

issue in honor of Albert Einstein on his 70th birthday.

Hevesy:1945:ERR

[Hev45] G. Hevesy. On the effect of Roentgen rays on cellular division. Reviews of Modern Physics, 17(2–3):102–111, April 1945. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.17.102;   http://rmp.aps.org/abstract/
RMP/v17/i2-3/p102_1.

Hirsh:1942:SXR

[Hir42] F. R. Hirsh, Jr. A summary of X-ray satellites. Reviews of Modern Physics, 14(1):45–54, January 1942. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
14.45;   http://rmp.aps.org/abstract/RMP/v14/i1/p45_1.

Hornyak:1948:ELL

[HL48] W. F. Hornyak and T. Lauritsen. Energy levels of light nuclei. Reviews of Modern Physics, 20(1):191–227, January 1948. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
20.191;   http://rmp.aps.org/abstract/RMP/v20/i1/p191_
1.

Herring:1949:TE

[HN49] Conyers Herring and M. H. Nichols. Thermionic emission. Reviews of Modern Physics, 21(2):185–270, April 1949. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.21.185;
 http://rmp.aps.org/abstract/RMP/v21/i2/p185_1.

Hoffmann:1949:KNR

[Hof49] Banesh Hoffmann. Kron’s non-Riemannian electrodynamics. Reviews of Modern Physics, 21(3):535–540, July 1949. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.21.535;
 http://rmp.aps.org/abstract/RMP/v21/i3/p535_1. Special

issue in honor of Albert Einstein on his 70th birthday.

Houston:1948:NVC

[Hou48] W. V. Houston. Normal vibrations of a crystal lattice. Reviews of Modern Physics, 20(1):161–165, January 1948. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.20.161;
 http://rmp.aps.org/abstract/RMP/v20/i1/p161_1.

Herzfeld:1942:CCP

[HS42] K. F. Herzfeld and A. L. Sklar. Color and constitution of polymethine dyes. Reviews of Modern Physics, 14(2–3):294–302, April 1942. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.14.294;   http://rmp.aps.org/abstract/
RMP/v14/i2-3/p294_1.

Hanson:1949:MNC

[HTW49] A. O. Hanson, R. F. Taschek, and J. H. Williams. Monoergic neutrons from charged particle reactions. Reviews of Modern Physics, 21(4):635–650, October 1949. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.21.635;
 http://rmp.aps.org/abstract/RMP/v21/i4/p635_1.

Hulthen:1945:CDM

[Hul45] Lamek Hulthén. Comments on the difficulties of the meson theory. Reviews of Modern Physics, 17(2–3):263–266, April 1945. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.17.263;   http://rmp.aps.org/abstract/
RMP/v17/i2-3/p263_1.

Heitler:1945:TCR

[HW45] W. Heitler and P. Walsh. Theory of cosmic-ray mesons. Reviews of Modern Physics, 17(2–3):252–262, April 1945. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.17.252;
 http://rmp.aps.org/abstract/RMP/v17/i2-3/p252_1.

Infeld:1949:MTP

[IS49] L. Infeld and A. Schild. On the motion of test particles in General Relativity. Reviews of Modern Physics, 21(3):408–413, July 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.21.408;   http://rmp.aps.org/abstract/

RMP/v21/i3/p408_1. Special issue in honor of Albert Einstein on his 70th birthday.

Jewett:1948:GNR

[Jew48] Frank B. Jewett. The genesis of the National Research Council and Millikan’s World War I work. Reviews of Modern Physics, 20(1):1–6, January 1948. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.20.1;   http://rmp.aps.
org/abstract/RMP/v20/i1/p1_1.

Jones:1946:SIT

[JF46] R. Clark Jones and W. H. Furry. The separation of isotopes by thermal diffusion. Reviews of Modern Physics, 18(2):151–224, April 1946. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.18.151;   http://rmp.aps.org/abstract/
RMP/v18/i2/p151_1.

Karapetoff:1944:STR

[Kar44] Vladimir Karapetoff. Special Theory of Relativity in hyperbolic functions. Reviews of Modern Physics, 16(1):33–52, January 1944. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.16.33;   http://rmp.aps.org/abstract/RMP/v16/
i1/p33_1.

Kauzmann:1942:DRC

[Kau42] Walter Kauzmann. Dielectric relaxation as a chemical rate process. Reviews of Modern Physics, 14(1):12–44, January 1942. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
14.12;   http://rmp.aps.org/abstract/RMP/v14/i1/p12_1.

Kittel:1949:PTF

[Kit49] Charles Kittel. Physical theory of ferromagnetic domains. Reviews of Modern Physics, 21(4):541–583, October 1949. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.21.541;
 http://rmp.aps.org/abstract/RMP/v21/i4/p541_1.

Klein:1945:SRQ

[KL45] O. Klein and J. Lindhard. Some remarks on the quantum theory of the superconductive state. Reviews of Modern Physics, 17(2–3):305–309, April 1945. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.17.305;   http://
rmp.aps.org/abstract/RMP/v17/i2-3/p305_1.

Klein:1949:TEF

[Kle49] O. Klein. On the thermodynamical equilibrium of fluids in gravitational fields. Reviews of Modern Physics, 21(3):531–533, July 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.21.531;   http://rmp.aps.org/abstract/

RMP/v21/i3/p531_1. Special issue in honor of Albert Einstein on his 70th birthday.

Kratz:1942:ASP

[KM42] H. R. Kratz and J. E. Mack. The absorption spectra of potassium, rubidium, and caesium. Reviews of Modern Physics, 14(2–3):104, April 1942. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.14.104;   http://rmp.aps.org/abstract/
RMP/v14/i2-3/p104_1.

Kellogg:1946:MBM

[KM46] J. B. M. Kellogg and S. Millman. The molecular beam magnetic resonance method. The radiofrequency spectra of atoms and molecules. Reviews of Modern Physics, 18(3):323–352, July 1946. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.18.323;   http://rmp.aps.org/abstract/
RMP/v18/i3/p323_1.

Konopinski:1943:BD

[Kon43] Emil Jan Konopinski. Beta-decay. Reviews of Modern Physics, 15(4):209–245, October 1943. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.15.209;   http://
rmp.aps.org/abstract/RMP/v15/i4/p209_1.

Korff:1942:OPC

[Kor42] S. A. Korff. The operation of proportional counters. Reviews of Modern Physics, 14(1):1–11, January 1942. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.14.1;
http://rmp.aps.org/abstract/RMP/v14/i1/p1_1.

Korff:1948:PND

[Kor48] Serge A. Korff. The production of nuclear disruptions by the cosmic radiation. Reviews of Modern Physics, 20(1):327–333, January 1948. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.20.327;   http://rmp.aps.org/abstract/
RMP/v20/i1/p327_1.

Korff:1949:CIM

[Kor49] Serge A. Korff. A critique of ionization measurements of nuclear disruptions produced by cosmic radiation. Reviews of Modern Physics, 21(1):75–81, January 1949. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
21.75;   http://rmp.aps.org/abstract/RMP/v21/i1/p75_1.

Kaufmann:1945:MGI

[KPC45] A. R. Kaufmann, S. T. Pan, and J. R. Clark. Magnetization of gold-iron and gold-nickel solid solutions. Reviews of Modern Physics, 17(1):87–92, January 1945. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
17.87;   http://rmp.aps.org/abstract/RMP/v17/i1/p87_1.

Kane:1949:ICR

[KSW49] Evan O. Kane, T. J. B. Shanley, and John A. Wheeler. Influence on the cosmic-ray spectrum of five heavenly bodies. Reviews of Modern Physics, 21(1):51–71, January 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
21.51;   http://rmp.aps.org/abstract/RMP/v21/i1/p51_1.

Lanczos:1949:LMR

[Lan49] Cornelius Lanczos. Lagrangian multiplier and Riemannian spaces. Reviews of Modern Physics, 21(3):497–502, July 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
21.497;   http://rmp.aps.org/abstract/RMP/v21/i3/p497_

1. Special issue in honor of Albert Einstein on his 70th birthday.

Latyshev:1947:IRM

[Lat47] G. D. Latyshev. The interaction of γ-rays with matter and the spectroscopy of γ-radiation. Reviews of Modern Physics, 19(2):132–145, April 1947. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.19.132;   http://rmp.aps.org/abstract/
RMP/v19/i2/p132_1.

Ladenburg:1949:LTB

[LB49] R. Ladenburg and D. Bershader. On laminar and turbulent boundary layer in supersonic flow. Reviews of Modern Physics, 21(3):510–515, July 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.21.510;   http://

rmp.aps.org/abstract/RMP/v21/i3/p510_1. Special issue in

honor of Albert Einstein on his 70th birthday.

Leighton:1948:VSS

[Lei48] Robert B. Leighton. The vibrational spectrum and specific heat of a face-centered cubic crystal. Reviews of Modern Physics, 20 (1):165–174, January 1948. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.20.165;   http://
rmp.aps.org/abstract/RMP/v20/i1/p165_1.

Lemaitre:1949:CAR

[Lem49] Georges Lemaître. Cosmological application of Relativity. Reviews of Modern Physics, 21(3):357–366, July 1949. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.21.357;
 http://rmp.aps.org/abstract/RMP/v21/i3/p357_1. Special

issue in honor of Albert Einstein on his 70th birthday.

Livingston:1946:ISC

[Liv46] M. Stanley Livingston. Ion sources for cyclotrons. Reviews of Modern Physics, 18(3):293–299, July 1946. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.18.293;
 http://rmp.aps.org/abstract/RMP/v18/i3/p293_1.

Loeb:1948:SFS

[Loe48] Leonard B. Loeb. Statistical factors in spark discharge mechanisms. Reviews of Modern Physics, 20(1):151–160, January 1948. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
20.151;   http://rmp.aps.org/abstract/RMP/v20/i1/p151_
1.

London:1945:PCL

[Lon45] F. London. Planck’s constant and low temperature transfer. Reviews of Modern Physics, 17(2–3):310–320, April 1945. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
17.310;   http://rmp.aps.org/abstract/RMP/v17/i2-3/p310_

Loofbourow:1940:BPB

[Loo40] John R. Loofbourow. Borderland problems in biology and physics. Reviews of Modern Physics, 12(4):267–358, October 1940. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
12.267;   http://rmp.aps.org/abstract/RMP/v12/i4/p267_
1.

Leprince-Ringuet:1949:PEE

[LR49] Louis Leprince-Ringuet. Photographic evidence for the existence of a very heavy meson. Reviews of Modern Physics, 21(1):42–43, January 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.21.42;   http://
rmp.aps.org/abstract/RMP/v21/i1/p42_1.

Livingood:1940:TIR

[LS40] J. J. Livingood and G. T. Seaborg. A table of induced radioactivities. Reviews of Modern Physics, 12(1):30–46, January 1940. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.12.30;   http://rmp.aps.org/abstract/RMP/v12/
i1/p30_1.

Margenau:1941:MEP

[Mar41] H. Margenau. Metaphysical elements in physics. Reviews of Modern Physics, 13(3):176–189, July 1941. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.13.176;
 http://rmp.aps.org/abstract/RMP/v13/i3/p176_1.

Marshak:1947:TSN

[Mar47] Robert E. Marshak. Theory of the slowing down of neutrons by elastic collision with atomic nuclei. Reviews of Modern Physics, 19(3):185–238, July 1947. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.19.185;   http://
rmp.aps.org/abstract/RMP/v19/i3/p185_1.

Manning:1940:ECR

[MB40] Millard F. Manning and Maurice E. Bell. Electrical conduction and related phenomena in solid dielectrics. Reviews of Modern Physics, 12(3):215–256, July 1940. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.12.215;
 http://rmp.aps.org/abstract/RMP/v12/i3/p215_1.

Mack:1942:ABA

[MB42] J. E. Mack and E. C. Barkofsky. Atomic beam apparatus for studying the atomic spectra of gases, especially hydrogen. Reviews of Modern Physics, 14(2–3):82–93, April 1942. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.14.82;
 http://rmp.aps.org/abstract/RMP/v14/i2-3/p82_1.

Morse:1944:SWRa

[MB44a] Philip M. Morse and Richard H. Bolt. Sound waves in rooms. Reviews of Modern Physics, 16(2):69–150, April 1944. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
16.69;   http://rmp.aps.org/abstract/RMP/v16/i2/p69_1.

Morse:1944:SWRb

[MB44b] Philip M. Morse and Richard H. Bolt. Sound waves in rooms. Reviews of Modern Physics, 16(3–4):324, July 1944. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.16.324;
 http://rmp.aps.org/abstract/RMP/v16/i3-4/p324_1.

Marinelli:1947:AEB

[MBH47] L. D. Marinelli, R. F. Brinckerhoff, and G. J. Hine. Average energy of beta-rays emitted by radioactive isotopes. Reviews of Modern Physics, 19(1):25–28, January 1947. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
19.25;   http://rmp.aps.org/abstract/RMP/v19/i1/p25_1.

McKellar:1942:IME

[McK42] Andrew McKellar. Intensity measurements on emission bands in cometary spectra. Reviews of Modern Physics, 14(2–3):179–189, April 1942. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.14.179;   http://rmp.aps.org/abstract/
RMP/v14/i2-3/p179_1.

Meggers:1942:ASR

[Meg42a] William F. Meggers. Atomic spectra of rare earth elements. Reviews of Modern Physics, 14(2–3):96–103, April 1942. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
14.96;   http://rmp.aps.org/abstract/RMP/v14/i2-3/p96_
1.

Meggers:1942:PSW

[Meg42b] William F. Meggers. The primary standard of wave-length. Reviews of Modern Physics, 14(2–3):59–63, April 1942. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.14.59;
 http://rmp.aps.org/abstract/RMP/v14/i2-3/p59_1.

Meissner:1942:AAB

[Mei42] Karl Wilh. Meissner. Application of atomic beams in spectroscopy. Reviews of Modern Physics, 14(2–3):68–78, April 1942. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
14.68;   http://rmp.aps.org/abstract/RMP/v14/i2-3/p68_
1.

Meitner:1945:ASS

[Mei45] Lise Meitner. An attempt to single out some fission processes of uranium by using the differences in their energy release. Reviews of Modern Physics, 17(2–3):287–291, April 1945. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.17.287;
 http://rmp.aps.org/abstract/RMP/v17/i2-3/p287_1.

Maze:1949:EPA

[MFDA49] R. Maze, A. Fréon, J. Daudin, and P. Auger. Extensive and penetrating atmospheric showers. Reviews of Modern Physics, 21(1):14–20, January 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.21.14;   http://
rmp.aps.org/abstract/RMP/v21/i1/p14_1.

Millikan:1949:AEH

[Mil49a] Robert A. Millikan. Albert Einstein on his seventieth birthday. Reviews of Modern Physics, 21(3):343–345, July 1949. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.21.343;
 http://rmp.aps.org/abstract/RMP/v21/i3/p343_1. Special

issue in honor of Albert Einstein on his 70th birthday.

Millikan:1949:PSE

[Mil49b] Robert A. Millikan. The present status of the evidence for the atom-annihilation hypothesis. Reviews of Modern Physics, 21 (1):1–13, January 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.21.1;   http://rmp.aps.
org/abstract/RMP/v21/i1/p1_1.

Mitchell:1948:UCC

[Mit48] Allan C. G. Mitchell. The use of coincidence counting methods in determining nuclear disintegration schemes. Reviews of Modern Physics, 20(1):296–304, January 1948. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.20.296;
 http://rmp.aps.org/abstract/RMP/v20/i1/p296_1.

Mayer:1942:CAS

[MM42] Maria Goeppert Mayer and Kenneth J. McCallum. Calculation of the absorption spectrum of Wurster’s Salts. Reviews of Modern Physics, 14(2–3):248–257, April 1942. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.14.248;
 http://rmp.aps.org/abstract/RMP/v14/i2-3/p248_1.

Morton:1946:EGT

[Mor46] G. A. Morton. Electron guns for television application. Reviews of Modern Physics, 18(3):362–378, July 1946. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.18.362;
 http://rmp.aps.org/abstract/RMP/v18/i3/p362_1.

Mulliken:1942:BIS

[MR42] Robert S. Mulliken and Carol A. Rieke. Bond integrals and spectra with an analysis of Kynch and Penney’s paper on the heat of sublimation of carbon. Reviews of Modern Physics, 14(2–3):259, April 1942. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.14.259;   http://rmp.aps.org/abstract/
RMP/v14/i2-3/p259_1.

Mrozowski:1942:ESI

[Mro42] S. Mrozowski. Emission spectrum of the ion CO_{2}^{+}. Reviews of Modern Physics, 14(2–3):216–218, April 1942. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.14.216;
 http://rmp.aps.org/abstract/RMP/v14/i2-3/p216_1.

Mrozowski:1944:FLL

[Mro44] S. Mrozowski. Forbidden lines in the laboratory. Reviews of Modern Physics, 16(3–4):153–174, July 1944. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.16.153;
 http://rmp.aps.org/abstract/RMP/v16/i3-4/p153_1.

Mulliken:1942:ESS

[Mul42a] Robert S. Mulliken. Electronic structures and spectra of triatomic oxide molecules. Reviews of Modern Physics, 14(2–3):204–215, April 1942. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.14.204;   http://rmp.aps.org/abstract/
RMP/v14/i2-3/p204_1.

Mulliken:1942:IC

[Mul42b] Robert S. Mulliken. Introduction to the conference. Reviews of Modern Physics, 14(2–3):57–58, April 1942. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.14.57;
 http://rmp.aps.org/abstract/RMP/v14/i2-3/p57_1.

Mulliken:1942:SUS

[Mul42c] Robert S. Mulliken. Structure and ultraviolet spectra of ethylene, butadiene, and their alkyl derivatives. Reviews of Modern Physics, 14(2–3):265–274, April 1942. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.14.265;   http:/
/rmp.aps.org/abstract/RMP/v14/i2-3/p265_1.

Nielsen:1944:IRA

[NN44] Alvin H. Nielsen and Harald H. Nielsen. The infra-red absorption spectrum of deutero-methyl chloride. Reviews of Modern Physics, 16(3–4):241–244, July 1944. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.16.241;   http:/
/rmp.aps.org/abstract/RMP/v16/i3-4/p241_1.

Neher:1948:CRE

[NR48] H. V. Neher and W. C. Roesch. Cosmic-ray effects from solar flares and magnetic storms. Reviews of Modern Physics, 20 (1):350–352, January 1948. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.20.350;   http://
rmp.aps.org/abstract/RMP/v20/i1/p350_1.

Nielsen:1944:ALG

[NTD44] J. Rud Nielsen, V. Thornton, and E. Brock Dale. The absorption laws for gases in the infra-red. Reviews of Modern Physics, 16(3–4):307–324, July 1944. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.16.307;   http://
rmp.aps.org/abstract/RMP/v16/i3-4/p307_1.

Newton:1949:LSE

[NW49] T. D. Newton and E. P. Wigner. Localized states for elementary systems. Reviews of Modern Physics, 21(3):400–406, July 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.21.400;   http://rmp.aps.org/abstract/

RMP/v21/i3/p400_1. Special issue in honor of Albert Einstein on his 70th birthday.

Oppenheimer:1949:CRC

[Opp49] J. R. Oppenheimer. Concluding remarks to cosmic-ray symposium. Reviews of Modern Physics, 21(1):181–183, January 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
21.181;   http://rmp.aps.org/abstract/RMP/v21/i1/p181_
1.

Oetjen:1944:IRSb

[OR44] R. A. Oetjen and H. M. Randall. The infra-red spectra of the isomeric octanes in the vapor phase. Reviews of Modern Physics, 16(3–4):265–270, July 1944. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.16.265;   http://
rmp.aps.org/abstract/RMP/v16/i3-4/p265_1.

Oetjen:1944:IRSa

[ORA44] R. A. Oetjen, H. M. Randall, and W. E. Anderson. The infrared spectra of the isomeric octanes in the liquid phase. Reviews of Modern Physics, 16(3–4):260–264, July 1944. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.16.260;
 http://rmp.aps.org/abstract/RMP/v16/i3-4/p260_1.

Pauli:1941:RFT

[Pau41] W. Pauli. Relativistic field theories of elementary particles. Reviews of Modern Physics, 13(3):203–232, July 1941. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.13.203;
 http://rmp.aps.org/abstract/RMP/v13/i3/p203_1.

Pauli:1943:DNM

[Pau43] W. Pauli. On Dirac’s new method of field quantization. Reviews of Modern Physics, 15(3):175–207, July 1943. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.15.175;
 http://rmp.aps.org/abstract/RMP/v15/i3/p175_1.

Pauli:1945:NBH

[Pau45] W. Pauli. Niels Bohr on his 60th birthday. Reviews of Modern Physics, 17(2–3):97–101, April 1945. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.17.97;   http://
rmp.aps.org/abstract/RMP/v17/i2-3/p97_1.

Pomerance:1942:IOG

[PB42] H. S. Pomerance and H. G. Beutler. Identification of orders and ghosts in grating spectra by diffracting slits. Reviews of Modern Physics, 14(2–3):66–67, April 1942. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.14.66;   http://
rmp.aps.org/abstract/RMP/v14/i2-3/p66_1.

Pauling:1948:RVE

[PE48] Linus Pauling and Fred J. Ewing. The ratio of valence electrons to atoms in metals and intermetallic compounds. Reviews of Modern Physics, 20(1):112–122, January 1948. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.20.112;
 http://rmp.aps.org/abstract/RMP/v20/i1/p112_1.

Pais:1949:NRP

[PE49] A. Pais and S. T. Epstein. Note on relativistic properties of self-energies. Reviews of Modern Physics, 21(3):445–446, July 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.21.445;   http://rmp.aps.org/abstract/

RMP/v21/i3/p445_1. Special issue in honor of Albert Einstein on his 70th birthday.

Pauli:1945:SCC

[PH45] W. Pauli and N. Hu. On the strong coupling case for spindependent interactions in scalar- and vector-pair theories. Reviews of Modern Physics, 17(2–3):267–286, April 1945. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.17.267;
 http://rmp.aps.org/abstract/RMP/v17/i2-3/p267_1.

Platt:1944:SOM

[PK44] J. R. Platt and H. B. Klevens. Spectroscopy of organic molecules in the vacuum ultraviolet. Reviews of Modern Physics, 16 (3–4):182–223, July 1944. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.16.182;   http://
rmp.aps.org/abstract/RMP/v16/i3-4/p182_1.

Pringsheim:1942:FPT

[Pri42] Peter Pringsheim. Fluorescence and phosphorescence of thalliumactivated potassium-halide phosphors. Reviews of Modern Physics, 14(2–3):132–138, April 1942. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.14.132;   http://
rmp.aps.org/abstract/RMP/v14/i2-3/p132_1.

Plesset:1948:DCF

[PS48a] Milton S. Plesset and Philip A. Shaffer, Jr. Drag in cavitating flow. Reviews of Modern Physics, 20(1):228–231, January 1948. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
20.228;   http://rmp.aps.org/abstract/RMP/v20/i1/p228_

Podolsky:1948:RGE

[PS48b] Boris Podolsky and Philip Schwed. Review of a generalized electrodynamics. Reviews of Modern Physics, 20(1): 40–50, January 1948. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.20.40;   http://
rmp.aps.org/abstract/RMP/v20/i1/p40_1.

Pauli:1949:IRR

[PV49] W. Pauli and F. Villars. On the invariant regularization in relativistic quantum theory. Reviews of Modern Physics, 21(3):434–444, July 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.21.434;   http://rmp.aps.org/abstract/

RMP/v21/i3/p434_1. Special issue in honor of Albert Einstein on his 70th birthday.

Potapenko:1948:NMD

[PW48] G. Potapenko and D. Wheeler, Jr. A new method of determining the relaxation time and the dipole moment of polar substances; its application to fatty acids. Reviews of Modern Physics, 20(1):143–150, January 1948. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.20.143;
 http://rmp.aps.org/abstract/RMP/v20/i1/p143_1.

Rabinowitch:1942:ETS

[Rab42] Eugene Rabinowitch. Electron transfer spectra and their photochemical effects. Reviews of Modern Physics, 14(2–3):112–131, April 1942. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.14.112;   http://rmp.aps.org/abstract/
RMP/v14/i2-3/p112_1.

Rabinowitch:1944:SPC

[Rab44] E. Rabinowitch. Spectra of porphyrins and chlorophyll. Reviews of Modern Physics, 16(3–4):226–235, July 1944. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.16.226;
 http://rmp.aps.org/abstract/RMP/v16/i3-4/p226_1.

Racah:1949:DTC

[Rac49] Giulio Racah. On the decomposition of tensors by contraction. Reviews of Modern Physics, 21(3):494–496, July 1949. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.21.494;
 http://rmp.aps.org/abstract/RMP/v21/i3/p494_1. Special

issue in honor of Albert Einstein on his 70th birthday.

Rossi:1949:DDN

[RAO^{+}49] B. Rossi, C. D. Anderson, J. R. Oppenheimer, G. E. Valley, and R. D. Sard. Discussion on the disintegration and nuclear absorption of mesons. Reviews of Modern Physics, 21 (1):31–37, January 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.21.31;   http://
rmp.aps.org/abstract/RMP/v21/i1/p31_1. Oppenheimer: Re-

mark on μ-Decay (pages 34–35).

Reynolds:1949:ELA

[Rey49] G. T. Reynolds. On the effect of lead above apparatus for the detection of low energy mesons. Reviews of Modern Physics, 21(1):122–123, January 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.21.122;   http://
rmp.aps.org/abstract/RMP/v21/i1/p122_1.

Rossi:1941:CRT

[RG41] Bruno Rossi and Kenneth Greisen. Cosmic-ray theory. Reviews of Modern Physics, 13(4):240–309, October 1941. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.13.240;
 http://rmp.aps.org/abstract/RMP/v13/i4/p240_1.

Robertson:1949:PVO

[Rob49] H. P. Robertson. Postulate versus observation in the Special Theory of Relativity. Reviews of Modern Physics, 21(3):378–382, July 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.21.378;   http://rmp.aps.org/abstract/

RMP/v21/i3/p378_1. Special issue in honor of Albert Einstein on his 70th birthday.

Rochester:1949:PPC

[Roc49] G. D. Rochester. The penetrating particles in cosmic-ray showers. Reviews of Modern Physics, 21(1):20–26, January 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
21.20;   http://rmp.aps.org/abstract/RMP/v21/i1/p20_1.

Roller:1948:MIU

[Rol48] Duane Roller. Millikan’s influence on undergraduate teaching. Reviews of Modern Physics, 20(1):26–30, January 1948. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
20.26;   http://rmp.aps.org/abstract/RMP/v20/i1/p26_1.

Rossi:1948:ICR

[Ros48] Bruno Rossi. Interpretation of cosmic-ray phenomena. Reviews of Modern Physics, 20(3):537–583, July 1948. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.20.537;
 http://rmp.aps.org/abstract/RMP/v20/i3/p537_1.

Rosen:1949:PRS

[Ros49a] Nathan Rosen. A particle at rest in a static gravitational field. Reviews of Modern Physics, 21(3):503–505, July 1949. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.21.503;
 http://rmp.aps.org/abstract/RMP/v21/i3/p503_1. Special

issue in honor of Albert Einstein on his 70th birthday.

Rossi:1949:EPC

[Ros49b] Bruno Rossi. Electrons and photons in cosmic rays. Reviews of Modern Physics, 21(1):104–112, January 1949. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.21.104;
 http://rmp.aps.org/abstract/RMP/v21/i1/p104_1.

Samuel:1946:DSC

[Sam46] R. Samuel. The dissociation spectra of covalent polyatomic molecules. Reviews of Modern Physics, 18(1):103–147, January 1946. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.18.103;   http://rmp.aps.org/abstract/
RMP/v18/i1/p103_1.

Schouten:1949:MFC

[Sch49] J. A. Schouten. On meson fields and conformal transformations. Reviews of Modern Physics, 21(3):421–424, July 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
21.421;   http://rmp.aps.org/abstract/RMP/v21/i3/p421_

1. Special issue in honor of Albert Einstein on his 70th birthday.

Seaborg:1944:TI

[Sea44] Glenn T. Seaborg. Table of isotopes. Reviews of Modern Physics, 16(1):1–32, January 1944. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.16.1;   http://rmp.aps.
org/abstract/RMP/v16/i1/p1_1.

Seitz:1946:CCA

[Sei46] Frederick Seitz. Color centers in alkali halide crystals. Reviews of Modern Physics, 18(3):384–408, July 1946. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.18.384;
 http://rmp.aps.org/abstract/RMP/v18/i3/p384_1.

Segre:1949:NI

[SH49] Emilio Segrè and A. C. Helmholz. Nuclear isomerism. Reviews of Modern Physics, 21(2):271–304, April 1949. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.21.271;
 http://rmp.aps.org/abstract/RMP/v21/i2/p271_1. See erra-

tum [SH50].

Segre:1950:ENI

[SH50] Emilio Segré and A. C. Helmholz. Erratum: Nuclear Isomerism. Reviews of Modern Physics, 22(3):237, July 1950. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
22.237;   http://rmp.aps.org/abstract/RMP/v22/i3/p237_

1. See [SH49].

Shapiro:1941:TNP

[Sha41] Maurice M. Shapiro. Tracks of nuclear particles in photographic emulsions. Reviews of Modern Physics, 13(1):

58–71, January 1941. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.13.58;   http://
rmp.aps.org/abstract/RMP/v13/i1/p58_1.

Shaffer:1944:DMV

[Sha44] Wave H. Shaffer. Degenerate modes of vibration and perturbations in polyatomic molecules. Reviews of Modern Physics, 16(3–4):245–259, July 1944. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.16.245;   http://
rmp.aps.org/abstract/RMP/v16/i3-4/p245_1.

Sheppard:1942:EEA

[She42a] S. E. Sheppard. The effects of environment and aggregation on the absorption spectra of dyes. Reviews of Modern Physics, 14(2–3):303–340, April 1942. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.14.303;   http:/
/rmp.aps.org/abstract/RMP/v14/i2-3/p303_1.  See errata

[She42b].

Sheppard:1942:EEE

[She42b] S. E. Sheppard. Errata: The Effects of Environment and Aggregation on the Absorption Spectra of Dyes. Reviews of Modern Physics, 14(4):410, October 1942. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.14.410;   http:/
/rmp.aps.org/abstract/RMP/v14/i4/p410_1. See [She42a].

Skellett:1946:BPR

[Ske46] A. M. Skellett. Beam production in radial beam tubes, beam power tubes, and other low voltage electronic devices. Reviews of Modern Physics, 18(3):379–383, July 1946. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.18.379;
 http://rmp.aps.org/abstract/RMP/v18/i3/p379_1.

Sklar:1942:EAS

[Skl42] A. L. Sklar. Electronic absorption spectra of benzene and its derivatives. Reviews of Modern Physics, 14(2–3):232–245, April 1942. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.14.232;   http://rmp.aps.org/abstract/
RMP/v14/i2-3/p232_1.

Slater:1946:ME

[Sla46] J. C. Slater. Microwave electronics. Reviews of Modern Physics, 18(4):441–512, October 1946. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.18.441;   http://
rmp.aps.org/abstract/RMP/v18/i4/p441_1.

Slater:1948:DLA

[Sla48] J. C. Slater. The design of linear accelerators. Reviews of Modern Physics, 20(3):473–518, July 1948. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.20.473;
 http://rmp.aps.org/abstract/RMP/v20/i3/p473_1.

Smyth:1945:AEM

[Smy45] H. D. Smyth. Atomic energy for military purposes. Reviews of Modern Physics, 17(4):351–471, October 1945. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.17.351;
 http://rmp.aps.org/abstract/RMP/v17/i4/p351_1.

Smythe:1948:ENG

[Smy48a] W. R. Smythe. Erratum: Narrow Gaps in Microwave Problems. Reviews of Modern Physics, 20(3):472, July 1948. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
20.472;   http://rmp.aps.org/abstract/RMP/v20/i3/p472_

1. See [Smy48b].

Smythe:1948:NGM

[Smy48b] W. R. Smythe. Narrow gaps in microwave problems. Reviews of Modern Physics, 20(1):175–180, January 1948. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.20.175;
 http://rmp.aps.org/abstract/RMP/v20/i1/p175_1. See erra-

tum [Smy48a].

Seaborg:1948:TI

[SP48] G. T. Seaborg and I. Perlman. Table of isotopes. Reviews of Modern Physics, 20(4):585–667, October 1948. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.20.585;
 http://rmp.aps.org/abstract/RMP/v20/i4/p585_1.

Sponer:1942:UAS

[Spo42] H. Sponer. Ultraviolet absorption spectra of substituted benzenes. Reviews of Modern Physics, 14(2–3):224–231, April 1942. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.14.224;   http://rmp.aps.org/abstract/
RMP/v14/i2-3/p224_1.

Sponer:1944:NUA

[Spo44] H. Sponer. The near ultraviolet absorption of pyridine vapor. Reviews of Modern Physics, 16(3–4):224–225, July 1944. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
16.224;   http://rmp.aps.org/abstract/RMP/v16/i3-4/p224_
1.

Sponer:1941:ESP

[ST41] H. Sponer and E. Teller. Electronic spectra of polyatomic molecules. Reviews of Modern Physics, 13(2):75–170, April 1941. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.13.75;   http://rmp.aps.org/abstract/RMP/v13/
i2/p75_1.

Stephens:1947:NHM

[Ste47] W. E. Stephens. The neutron-hydrogen mass difference and the neutron mass. Reviews of Modern Physics, 19(1):19–24, January 1947. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.19.19;   http://rmp.aps.org/abstract/RMP/v19/
i1/p19_1.

Stern:1949:MHI

[Ste49a] Martin O. Stern. The masses of the heavy isotopes. Reviews of Modern Physics, 21(2):316–321, April 1949. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.21.316;
 http://rmp.aps.org/abstract/RMP/v21/i2/p316_1.

Stern:1949:TE

[Ste49b] Otto Stern. On the term k ln n! in the entropy. Reviews of Modern Physics, 21(3):534–535, July 1949. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.21.534;
 http://rmp.aps.org/abstract/RMP/v21/i3/p534_1. Special

issue in honor of Albert Einstein on his 70th birthday.

Struve:1944:RPI

[Str44] Otto Struve. Recent progress in the interpretation of stellar spectra. Reviews of Modern Physics, 16(3–4):286–300, July 1944. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.16.286;   http://rmp.aps.org/abstract/
RMP/v16/i3-4/p286_1.

Strong:1948:PMN

[Str48] John Strong. The parastatic moving needle galvanometer. Reviews of Modern Physics, 20(1):180–184, January 1948. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.20.180;
 http://rmp.aps.org/abstract/RMP/v20/i1/p180_1.

Straus:1949:SRE

[Str49] E. G. Straus. Some results in Einstein’s Unified Field Theory. Reviews of Modern Physics, 21(3):414–420, July 1949. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.21.414;
 http://rmp.aps.org/abstract/RMP/v21/i3/p414_1. Special

issue in honor of Albert Einstein on his 70th birthday.

Swann:1941:RTE

[Swa41a] W. F. G. Swann. The relation of theory to experiment in physics. Reviews of Modern Physics, 13(3):190–196, July 1941. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
13.190;   http://rmp.aps.org/abstract/RMP/v13/i3/p190_

Swann:1941:RFL

[Swa41b] W. F. G. Swann. Relativity, the Fitzgerald–Lorentz contraction, and quantum theory. Reviews of Modern Physics, 13(3):197–202, July 1941. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.13.197;   http://rmp.aps.org/abstract/
RMP/v13/i3/p197_1.

Swings:1942:MBC

[Swi42] P. Swings. Molecular bands in cometary spectra. Identifications. Reviews of Modern Physics, 14(2–3):190–194, April 1942. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.14.190;   http://rmp.aps.org/abstract/
RMP/v14/i2-3/p190_1.

Sigurgeirsson:1949:EEP

[SY49] Thorbjorn Sigurgeirsson and K. Alan Yamakawa. Electron emitting power of stopped mesons. Reviews of Modern Physics, 21 (1):124–132, January 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.21.124;   http://
rmp.aps.org/abstract/RMP/v21/i1/p124_1.

Taub:1949:SMS

[Tau49] A. H. Taub. A special method for solving the Dirac equations. Reviews of Modern Physics, 21(3):388–392, July 1949. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.21.388;
 http://rmp.aps.org/abstract/RMP/v21/i3/p388_1. Special

issue in honor of Albert Einstein on his 70th birthday.

Turkevich:1942:RSN

[TF42] Anthony Turkevich and Mark Fred. Rotational structure of the near ultraviolet bands of benzene. Reviews of Modern Physics, 14(2–3):246–247, April 1942. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.14.246;   http://
rmp.aps.org/abstract/RMP/v14/i2-3/p246_1.

Tolman:1948:IPE

[TF48] Richard C. Tolman and Paul C. Fine. On the irreversible production of entropy. Reviews of Modern Physics, 20(1):

51–77, January 1948. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.20.51;   http://
rmp.aps.org/abstract/RMP/v20/i1/p51_1.

Project:1946:NFF

[The46] The Plutonium Project. Nuclei formed in fission: Decay characteristics, fission yields, and chain relationships. Reviews of Modern Physics, 18(4):513–544, October 1946. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.18.513;
 http://rmp.aps.org/abstract/RMP/v18/i4/p513_1. See erra-

tum [Ano47].

Thomas:1945:RI

[Tho45] L. H. Thomas. Relativistic invariance. Reviews of Modern Physics, 17(2–3):182–186, April 1945. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.17.182;   http://
rmp.aps.org/abstract/RMP/v17/i2-3/p182_1.

Turner:1940:NF

[Tur40] Louis A. Turner. Nuclear fission. Reviews of Modern Physics, 12(1):1–29, January 1940. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.12.1;   http://rmp.aps.
org/abstract/RMP/v12/i1/p1_1. This paper surveys the early

work from 1934 to 1939 on nuclear fission, including more than 100 from 1939 alone following the news of the Hahn and Strassmann work that Niels Bohr brought to the USA in January 1939.

Turner:1945:MHN

[Tur45] Louis A. Turner. The missing heavy nuclei. Reviews of Modern Physics, 17(2–3):292–296, April 1945. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.17.292;
 http://rmp.aps.org/abstract/RMP/v17/i2-3/p292_1.

Tiomno:1949:CER

[TW49a] J. Tiomno and John A. Wheeler. Charge-exchange reaction of the μ-meson with the nucleus. Reviews of Modern Physics, 21 (1):153–165, January 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.21.153;   http://
rmp.aps.org/abstract/RMP/v21/i1/p153_1.

Tiomno:1949:ESE

[TW49b] J. Tiomno and John A. Wheeler. Energy spectrum of electrons from meson decay. Reviews of Modern Physics, 21(1):144–152, January 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.21.144;   http://rmp.aps.org/abstract/
RMP/v21/i1/p144_1.

Vallarta:1949:GRE

[Val49] M. S. Vallarta. Galactic rotation effect and the origin of cosmic radiation. Reviews of Modern Physics, 21(3):356, July 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
21.356;   http://rmp.aps.org/abstract/RMP/v21/i3/p356_

1. Special issue in honor of Albert Einstein on his 70th birthday.

VanBiesbroeck:1942:STC

[Van42] G. Van Biesbroeck. Structural texture of comets. Reviews of Modern Physics, 14(2–3):160–163, April 1942. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.14.160;
 http://rmp.aps.org/abstract/RMP/v14/i2-3/p160_1.

VanVleck:1945:STF

[Van45] J. H. Van Vleck. A survey of the theory of ferromagnetism. Reviews of Modern Physics, 17(1):27–47, January 1945. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
17.27;   http://rmp.aps.org/abstract/RMP/v17/i1/p27_1.

vonKarman:1949:CST

[vKL49] Theodore von Kármán and C. C. Lin. On the concept of similiarity in the theory of isotropic turbulence. Reviews of Modern Physics, 21(3):516–519, July 1949. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.21.516;
 http://rmp.aps.org/abstract/RMP/v21/i3/p516_1. Special

issue in honor of Albert Einstein on his 70th birthday.

vonLaue:1949:AET

[vL49] M. von Laue. Zu Albert Einsteins 70-tem Geburtstag. Reviews of Modern Physics, 21(3):348–349, July 1949. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.21.348;
 http://rmp.aps.org/abstract/RMP/v21/i3/p348_1. Special

issue in honor of Albert Einstein on his 70th birthday.

VanVleck:1945:SCB

[VW45] J. H. Van Vleck and V. F. Weisskopf. On the shape of collisionbroadened lines. Reviews of Modern Physics, 17(2–3):227–236, April 1945. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.17.227;   http://rmp.aps.org/abstract/
RMP/v17/i2-3/p227_1.

Wannier:1945:SPC

[Wan45] G. H. Wannier. The statistical problem in cooperative phenomena. Reviews of Modern Physics, 17(1):50–60, January 1945. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
17.50;   http://rmp.aps.org/abstract/RMP/v17/i1/p50_1.

Weaver:1948:SFW

[Wea48] Warren Weaver. Statistical freedom of the will. Reviews of Modern Physics, 20(1):31–34, January 1948. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
20.31;   http://rmp.aps.org/abstract/RMP/v20/i1/p31_1.

Weisskopf:1949:RDT

[Wei49] Victor F. Weisskopf. Recent developments in the theory of the electron. Reviews of Modern Physics, 21(2):305–315, April 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.21.305;   http://rmp.aps.org/abstract/
RMP/v21/i2/p305_1.

Wentzel:1947:RRM

[Wen47] G. Wentzel. Recent research in meson theory. Reviews of Modern Physics, 19(1):1–18, January 1947. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.19.1;
http://rmp.aps.org/abstract/RMP/v19/i1/p1_1.

Wheeler:1945:IAM

[WF45] John Archibald Wheeler and Richard Phillips Feynman. Interaction with the absorber as the mechanism of radiation. Reviews of Modern Physics, 17(2–3):157–181, April 1945. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.17.157;
 http://rmp.aps.org/abstract/RMP/v17/i2-3/p157_1.

Wheeler:1949:CET

[WF49] John Archibald Wheeler and Richard Phillips Feynman. Classical electrodynamics in terms of direct interparticle action. Reviews of Modern Physics, 21(3):425–433, July 1949. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.21.425;
 http://rmp.aps.org/abstract/RMP/v21/i3/p425_1. Special

issue in honor of Albert Einstein on his 70th birthday.

Wheeler:1949:SCE

[Whe49] John A. Wheeler. Some consequences of the electromagnetic interaction between μ^{−}-mesons and nuclei. Reviews of Modern Physics, 21(1):133–143, January 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.21.133;   http://
rmp.aps.org/abstract/RMP/v21/i1/p133_1.

Whipple:1942:MEU

[Whi42] Fred L. Whipple. Meteors and the Earth’s upper atmosphere. Reviews of Modern Physics, 14(2–3):139, April 1942. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.14.139;
 http://rmp.aps.org/abstract/RMP/v14/i2-3/p139_1.

Whipple:1943:MEU

[Whi43] Fred L. Whipple. Meteors and the Earth’s upper atmosphere. Reviews of Modern Physics, 15(4):246–264, October 1943. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
15.246;   http://rmp.aps.org/abstract/RMP/v15/i4/p246_

Wildt:1942:GAC

[Wil42a] Rupert Wildt. The geochemistry of the atmosphere and the constitution of the terrestrial planets. Reviews of Modern Physics, 14(2–3):151–159, April 1942. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.14.151;   http://
rmp.aps.org/abstract/RMP/v14/i2-3/p151_1.

Williams:1942:AOE

[Wil42b] W. Ewart Williams. Attempts at obtaining excitation of an atomic beam of monatomic hydrogen. Reviews of Modern Physics, 14(2–3):94–95, April 1942. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.14.94;   http://
rmp.aps.org/abstract/RMP/v14/i2-3/p94_1.

Williams:1942:CUR

[Wil42c] W. Ewart Williams. Construction and use of reflection echelons. Reviews of Modern Physics, 14(2–3):64, April 1942. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.14.64;
 http://rmp.aps.org/abstract/RMP/v14/i2-3/p64_1.

Williams:1945:AOS

[Wil45] E. J. Williams. Application of ordinary space-time concepts in collision problems and relation of classical theory to Born’s approximation. Reviews of Modern Physics, 17(2–3):217–226, April 1945. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.17.217;   http://rmp.aps.org/abstract/
RMP/v17/i2-3/p217_1.

Winans:1944:PSR

[Win44] J. Gibson Winans. Partial selection rule for sensitized fluorescence. Reviews of Modern Physics, 16(3–4):175–181, July 1944. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.16.175;   http://rmp.aps.org/abstract/
RMP/v16/i3-4/p175_1.

Wang:1945:TBM

[WU45] Ming Chen Wang and G. E. Uhlenbeck. On the theory of the Brownian motion II. Reviews of Modern Physics, 17(2–3):323–342, April 1945. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.17.323;   http://rmp.aps.org/abstract/
RMP/v17/i2-3/p323_1.

Yukawa:1949:MMM

[Yuk49] Hideki Yukawa. Models and methods in the meson theory. Reviews of Modern Physics, 21(3):474–479, July 1949. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.21.474;
 http://rmp.aps.org/abstract/RMP/v21/i3/p474_1. Special

issue in honor of Albert Einstein on his 70th birthday.

Zener:1945:FSS

[Zen45] Clarence Zener. The fracture stress of steel. Reviews of Modern Physics, 17(1):20–26, January 1945. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
17.20;   http://rmp.aps.org/abstract/RMP/v17/i1/p20_1.

Zwicky:1940:TN

[Zwi40] F. Zwicky. Types of novae. Reviews of Modern Physics, 12 (1):66–85, January 1940. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.12.66;   http://
rmp.aps.org/abstract/RMP/v12/i1/p66_1.